### v2

# Asynchronous Federated Stochastic Optimization for Heterogeneous Objectives Under Arbitrary Delays

Abstract
Federated learning (FL) was recently proposed to securely train models with data held over multiple locations (“clients”) under the coordination of a central server. Two major challenges hindering the performance of FL algorithms are long training times caused by straggling clients, and a decline in model accuracy under non-iid local data distributions (“client drift”). In this work, we propose and analyze AsynchRonous Exact Averaging (AREA), a new stochastic (sub)gradient algorithm that utilizes asynchronous communication to speed up convergence and enhance scalability, and employs client memory to correct the client drift caused by variations in client update frequencies. Moreover, AREA is, to the best of our knowledge, the first method that is guaranteed to converge under arbitrarily long delays, without the use of delay-adaptive stepsizes, and (i) for strongly convex, smooth functions, asymptotically converges to an error neighborhood whose size depends only on the variance of the stochastic gradients used with respect to the number of iterations, and (ii) for convex, non-smooth functions, matches the convergence rate of the centralized stochastic subgradient method up to a constant factor, which depends on the average of the individual client update frequencies instead of their minimum (or maximum). Our numerical results validate our theoretical analysis and indicate AREA outperforms state-of-the-art methods when local data are highly non-iid, especially as the number of clients grows.
1 Introduction
Federated learning (FL) is a distributed machine learning paradigm that trains a shared model across multiple decentralized data sources, without the communication of data points [21, 29]. A common formulation of FL involves a finite-sum optimization problem of the form
$\min_{x\in\mathcal{X}}\quad f(x)=\sum_{i=1}^{n}f_{i}(x),$ (1)
where the $i^{th}$ data-owning source (client $i\in\{1,...,n\}$) has private access to the function $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ representing the loss over its local data, $x\in\mathbb{R}^{p}$ is the vector of model parameters, and $\mathcal{X}\subseteq\mathbb{R}^{p}$ is a constraint set. Throughout this work, we will refer to $f(\cdot)$ and the functions $f_{i}(\cdot)$ for $i\in\{1,...,n\}$ as the global and local loss functions, respectively.
At each round of a typical FL algorithm, clients use their local data to independently optimize the model parameters and share their estimates with the server. The server then produces the global model via an aggregation process (e.g., weighted averaging). Although FL enables learning rich models from decentralized, private data that may otherwise not be available, it may suffer from long training times due to the presence of slow clients (“stragglers”). Asynchronous communication can speed up convergence in the presence of stragglers by preventing fast clients from remaining idle, in addition to removing the need for synchronization which may be infeasible in large networks.
Asynchronous iterations, however, are difficult to analyze theoretically, as evidenced from the various frameworks proposed for this purpose over the years [27, 16, 15, 14]. The assumption of bounded client delays is frequently adopted to guarantee convergence (e.g., [26, 3, 44, 20, 34]), a setting known in the literature as “partially asynchronous” [4]; however, this assumption is largely impractical in FL, where clients behave inconsistenly over time. A more realistic model of asynchronous communication over networks is “total asynchrony” [4], i.e. arbitrarily long delays between iterations. In the FL setting, totally asynchronous iterations are typically analyzed under delay-adaptive stepsizes [40, 20, 51, 30, 15, 51], which may require small learning rates to accommodate long delays and hence may result in slow convergence. Unlike existing approaches, we adopt both unbounded delays and delay-independent stepsizes, and recover explicit convergence rates with respect to the number of iterations for non-iid local data distributions and (i) strongly convex and smooth functions and (ii) convex and non-smooth functions. To the best of our knowledge, the only work providing convergence guarantees in the non-iid data setting, under strong convexity and smoothness and linearly growing, unbounded delays is [17]; however, the resulting rates grow with the maximum delay between iterations. Conversely, we adopt a fully probabilistic framework (i.e., we make assumptions only on the statistics of client delays) and recover rates depending only on the average client update frequency, rendering our method less sensitive to fast or slow clients. In the case of convex and non-smooth functions and non-iid local data, the algorithmic framework in [14] does not impose restrictions on the delays or the stepsize; however, the resulting rates depend on the maximum client sampling probability. In contrast to [14], we allow clients to update the server at their own pace, and recover rates depending only on the average probabilities.
Another challenge hampering the performance of FL, addressed in this work, is the phenomenon of “client drift” [19, 48]. To reduce communication overhead, the seminal FL algorithm FedAvg [29], also known as local stochastic gradient descent (SGD), executes multiple consecutive SGD steps at each client before aggregations at the server. Under non-iid local data, however, performing multiple SGD steps may cause clients to converge close to the minima of their local losses instead of the minimum of the global loss, degrading the quality of the aggregated model (client drift) [24]. Asynchronous communication may exacerbate this deviation, with frequently updating clients being over-represented and slow clients underrepresented in the global model. Existing results for asynchronous FL (e.g., [54, 20, 46]) indicate that the convergence error grows with the product of client delays and the heterogeneity between the local datasets, even when a single SGD step is executed locally. To eliminate the client drift caused by variations in client update frequencies, we propose an error compensation scheme that utilizes client memory for the storage of correction terms. The effectiveness of our approach is validated by our theoretical results, which are independent of the heterogeneity between the local datasets when one SGD step is executed locally. A number of existing methods employ similar mechanisms for the same purpose [3, 17, 50, 53], but unlike this work, they do not confirm the elimination of client drift theoretically.
Contributions. We propose and analyze AsynchRonous Exact Averaging (AREA), a new stochastic (sub)gradient method that addresses the challenges of stragglers and scalability in FL by adopting asynchronous communication. To correct the client drift caused by heterogeneous client update frequencies, AREA utilizes a memory-based error compensation scheme. Moreover, our theoretical results do not require the unrealistic assumption of bounded delays to guarantee convergence, making AREA a totally asynchronous method. Our contributions can be summarized as follows:
  • We propose AREA, a new stochastic (sub)gradient method for FL with non-iid data, which utilizes asynchronous updates to address the issues of limited scalability and slow convergence caused by straggling clients.
  • AREA employs a memory-based compensation mechanism to correct the client drift caused by heterogeneous client update frequencies. Although other works utilize similar schemes [3, 17, 50, 53], we are the first to provide tight bounds which theoretically confirm the elimination of client drift for strongly convex and smooth functions. For convex and non-smooth functions and non-iid data, we are the first to report convergence rates matching the centralized projected stochastic subgradient method [7], up to a constant factor that depends only on the average client update frequency. Conversely, the only known rate in this setting grows with the maximum client update frequency [14], which may be significantly larger. Moreover, [14] adopts a biased sampling scheme for handling non-iid data, while we allow clients to update the server at their own pace. For a detailed comparison between our results and the state-of-the-art in the settings we study, see Tables 1 and 2 and Section 4.1.
  • Our analysis does not require bounded client delays to guarantee convergence, making AREA a totally asynchronous algorithm [4], without the use of delay-adaptive stepsizes.
  • For our theoretical analysis we prove an extended version (Lemma 20 in Appendix A.1) of the well-known result of Robbins–Siegmund (see [38], Lemma 11, Ch. 2), which is often used to derive the almost sure convergence of random sequences. This extended result can be used to construct suitable Lyapunov functions for tracking the progress of asynchronous iterations, and could be of independent interest in the probabilistic study of algorithms.
[…]
Table 1: Summary of asynchronous FL methods analyzed under strong convexity and smoothness. We refer to the asynchronous version of SGD as AsyncSGD. The works [43, 30, 15] adopt delay-adaptive stepsizes, while [17] assumes unbounded delays restricted to linear growth. For all methods, we consider the case where a single SGD is perfomed locally at each client before communicating with the server. We note that, although [3, 17] may correct the client drift due to asynchrony in practice, this is not verified theoretically; see Section 4.1 for more details.
Reference Non-iid data Unbounded delays No client drift
AsyncSGD [30] N/A
AsyncSGD [15] N/A
AsyncSGD [43] N/A
AsyncSGD [8]
AsyncCommSGD [3]
MIFA [17]
AREA (this work)
Table 2: Summary of asynchronous FL methods analyzed under convex and non-smooth functions $f_{i}(\cdot)$ in Problem (1). We refer to the asynchronous version of SGD as AsyncSGD.
Reference Non-iid data No client sampling
AsyncSGD [30]
AsyncSGD [14]
AREA (this work)
2 Related Work
Client drift in FL. Several methods have been proposed for tackling the client drift due to multiple local SGD steps in synchronous FL [23, 36, 19, 48]; however, none eliminate it completely [32]. FedLin [32] and Scaffnew [31] use gradient correcting terms to steer convergence towards the correct minimum, but require synchronous communication which limits their practical applicability. A small subset of asynchronous FL methods employ memory to remedy the client drift due to heterogeneous client update frequencies [3, 17, 50, 53]; however, they all require bounded client delays, with the exception of [17] which adopts linearly growing unbounded delays.
Asynchronous FL. Existing approaches for addressing stale updates and heterogeneous client update frequencies in asynchronous FL include (i) semi-synchronous strategies, such as adjusting the workload at each client [8, 25], or forcing clients to communicate inexact updates upon server request [55, 22] and (ii) fully asynchronous techniques, such as custom aggregation policies [35, 16, 52, 49] and delay adaptive stepsizes [40, 10, 20, 51, 30]. However, all methods listed above suffer from client drift due to variations in client update frequencies under non-iid data. Moreover, with the exception of the works that employ delay-adaptive stepsizes, all aforementioned works require bounded client delays to guarantee convergence. As we previously noted, methods utilizing delay-adaptive stepsizes may adopt small learning rates to compensate for extended delays which may slow down convergence.
AsynchronousSGD. Asynchronous SGD and its variants are the cornerstones of large-scale parallel and distributed optimization [47, 11, 39, 37, 45]. The convergence properties of parallel and distributed SGD [13, 1, 20, 43, 56, 26, 41, 2], and local SGD [42, 8, 30] with asynchronous or delayed updates have been studied extensively in the literature. For a detailed comparison between our results and existing results on asynchronous SGD in the settings we study see Tables 1 and 2.
3 AREA: An Asynchronous Method for Federated Stochastic Optimization
We propose a new asynchronous stochastic (sub)gradient algorithm, AREA, to solve Problem (1) in a distributed manner under arbitrary delays. We assume that client $i\in[n]$ can obtain only a stochastic approximation of the true (sub)gradient of $f_{i}(\cdot)$ (e.g., stochatic (sub)gradient, mini-batch), satisfying the following standard assumption.
Assumption 1.
Let $d_{i}(x)$ be the (sub)gradient of $f_{i}(\cdot)$ at $x\in\mathbb{R}^{p}$, and let $g_{i}(x)$ be a stochastic approximation of $d_{i}(x)$. Then for all $i\in[n]$ and some $\sigma>0$, we have: $\mathbb{E}[g_{i}(x)-d_{i}(x)|x]=0$ and $\mathbb{E}[\|g_{i}(x)-d_{i}(x)\|^{2}|x]\leq\sigma^{2}$, where the expectation is taken over all possible realizations of $g_{i}(x)$.
Algorithm 1 AREA (server)
Initialization: initialize model $x_{s}\in\mathcal{X}\subseteq\mathbb{R}^{p}$, aggregator $u_{s}=0\in\mathbb{R}^{p}$, stepsize $\alpha>0$ and stepsize update function $h:\mathbb{R}\rightarrow\mathbb{R}$; broadcast $x_{s}$ and $\alpha$ to all clients $i\in[n]$
while convergence is not reached do
     if a request from client $i$ is received then
         receive $m_{i}$ from client $i$ and update $u_{s}\leftarrow u_{s}+\frac{1}{n}m_{i}$
         update $\alpha\leftarrow h(\alpha)$ and send $x_{s}$ and $\alpha$ to client $i$
     if the aggregation criterion is satisfied then
         update $x_{s}\leftarrow x_{s}+u_{s}$, $u_{s}\leftarrow 0$
send stopping signal to all clients
The procedure of AREA at the server is summarized in Algorithm 1. The server maintains the global model $x_{s}\in\mathcal{X}$, an aggregator $u_{s}\in\mathbb{R}^{p}$, and a stepsize update function $h:\mathbb{R}\rightarrow\mathbb{R}$ (e.g., $h(x)=x$ for constant stepsize). At initialization, the server sets $u_{s}=0\in\mathbb{R}^{p}$, and initializes $x_{s}\in\mathcal{X}$ and the stepsize $\alpha$ and broadcasts them to all clients. Whenever an update $m_{i}\in\mathbb{R}^{p}$ arrives from some client $i\in[n]$, the server adds $m_{i}/n$ to $u_{s}$, constructs the new stepsize $h(\alpha)$ and sends it to client $i$ with $x_{s}$. In addition, the server is triggered whenever a pre-defined aggregation criterion is satisfied; the server then updates $x_{s}$ by adding to it the contents of $u_{s}$ and resets the value of $u_{s}$ to zero. These steps are repeated until convergence is reached and the server broadcasts a stopping signal to all clients.
Remark 2.
The aggregation criterion can be selected on a case-by-case basis to meet application demands. Possible options include aggregating every $\Delta\in\mathbb{N}$ client communications (e.g., [35]), after fixed time intervals (e.g., [16]), or even randomly. We note that it is possible to incorporate secure aggregation (SecAgg) to AREA for protection against honest-but-curious server attacks [6, 35].
Algorithm 2 AREA (client $i$)
Initialization: receive global model $x_{s}$ and set $y_{i}=x_{i}=x_{s}$
while stopping signal is not received do
     receive $x_{s}$ and $\alpha$ from server and update $x_{i}\leftarrow P_{\mathcal{X}}[x_{s}-\alpha g_{i}(x_{s})]$
     send $m_{i}=x_{i}-y_{i}$ to server and request $x_{s}$
     update $y_{i}\leftarrow x_{i}$
The process of AREA at client $i\in[n]$ is shown in Algorithm 2. Client $i\in[n]$ maintains two variables $x_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{X}$, storing the two most recent local estimates of the global model. During the initialization phase, clients receive $\alpha$ and $x_{s}$ and set $y_{i}=x_{i}=x_{s}$. The following steps are then repeated by all clients until a stopping signal from the server is received: whenever a global model $x_{s}$ and a stepsize value $\alpha$ are received from the server, client $i\in[n]$ computes at its own pace a stochastic (sub)gradient approximation $g_{i}(x_{s})$ of $f_{i}(\cdot)$ and takes a stochastic (sub)gradient step of length $\alpha_{k}$. The result of this operation is projected to $\mathcal{X}$ and stored in $x_{i}$. Next, client $i$ constructs the message $m_{i}=x_{i}-y_{i}$ and sends it to the server along with a request for the most recent version of $x_{s}$. Once client $i$ has communicated $m_{i}$ to the server, it stores the current value of $x_{i}$ in $y_{i}$.
Remark 3.
The communication round-based FedLaAvg (Algorithm 2 in [53]) with a single local SGD step at each client follows a similar procedure to AREA; however, FedLaAvg utilizes a stricter aggregation policy, where only the updates from clients who have not recently communicated with the server are used in the aggregation and the rest are discarded.
4 Convergence analysis
We now present our theoretical results on the convergence of AREA. We first set up our probabilistic framework for deriving the iterates of AREA and list our assumptions, then conclude with our theorems. The proofs for all theorems and intermediate lemmas can be found in Appendix A.1.
To capture more accurately the inconsistent behavior of clients in the FL setting, we adopt a probabilistic setup that allows for arbitrarily long delays between updates. Specifically, we define as an iteration of AREA every occurrence of one of the events $E_{i}$ for $i\in[n]$ and $E_{s}$, defined as follows: $E_{i}\triangleq\{\text{The server receives a request by client $i$}\}$ for $i\in[n]$, and $E_{s}\triangleq\{\text{The aggregation criterion is satisfied}\}$. We assume that the events $E_{s}$ and $E_{i}$ for $i\in[n]$ follow independent stochastic processes, namely, the time elapsed between two consecutive updates from client $i$ is independent from the time between two updates from client $j\neq i$ and from the time between two aggregations. [^1] This assumption does not cover aggregation schemes where the receipt of a fixed number of client updates triggers a global update (e.g., [35]); however, since our theoretical results are derived under the assumption of no information on the arrivals of client updates, they can be interpreted as a worst case/upper bounds for aggregation strategies making use of this information. Let $J_{k}\in\{1,2,...,n,s\}$ be a random variable describing the index $j$ of the event $E_{j}$ occurring at iteration $k$. We define $p_{i}\triangleq\mathbb{P}(J_{k}=i)>0$ for $i\in[n]$ and $p_{s}\triangleq\mathbb{P}(J_{k}=s)>0$.
Example 4.
A simple instance accommodated by our setup is modeling the time $t_{i}>0$ between two updates from client $i\in[n]$ and the time $t_{s}>0$ between two aggregations as independent exponential random variables with different parameters, i.e., $t_{i}\sim\exp(\lambda_{i})$ for $i\in[n]$ and $t_{s}\sim\exp(\lambda_{s})$. Then $\mathbb{P}(J_{k}=j)=\mathbb{P}(\min\{t_{1},...,t_{n},t_{s}\}=t_{j})$ which yields $p_{i}=\lambda_{i}/\bar{\lambda}$ for $i\in[n]$ and $p_{s}=\lambda_{s}/\bar{\lambda}$, where $\bar{\lambda}=\lambda_{s}+\sum_{i=1}^{n}\lambda_{i}$. This framework, sometimes referred to as Poisson clocks, is widely used to model asynchronous communication over networks, e.g., [45, 28, 9].
Unlike concurrent works, we do not impose deterministic bounds on the maximum or average client delays, or the growth of the delay sequence, and allow for heterogeneous client update frequencies. Moreover, modeling the satisfaction of the aggregation criterion as a random variable permits the adoption of a wide variety of stochastic and deterministic aggregation policies. [^2] Deterministic aggregation policies can be considered as special cases of the stochastic ones.
Let us define the Bernoulli random variables $\phi_{s,k}$ (resp. $\phi_{i,k}$) taking the value $1$ if $J_{k}=s$ (resp. $J_{k}=i$) at the $k^{th}$ iteration and 00 otherwise; then $\mathbb{E}[\phi_{s,k}|\mathcal{F}_{k-1}]=p_{s}$ and $\mathbb{E}[\phi_{i,k}|\mathcal{F}_{k-1}]=p_{i}$ for $i\in[n]$ and $k\in\mathbb{N}$. Moreover, assuming that the server keeps a count $k\in\mathbb{N}_{0}$ of the events $E_{i}$ for $i\in[n]$ and $E_{s}$, define $\alpha_{k}=h(\alpha_{k-1})$ for $k\in\mathbb{N}$ (we note that this counter is not necessary in practice due to the use of constant stepsizes; we only employ it for theoretical purposes). The iterates of AREA for $k\in\mathbb{N}_{0}$ can be expressed as
$x_{i,k+1}$ $=x_{i,k}+\phi_{i,k+1}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\right),\quad\forall i\in[n],$
$y_{i,k+1}$ $=y_{i,k}+\phi_{i,k+1}(x_{i,k}-y_{i,k}),\quad\forall i\in[n],$
$u_{s,k+1}$ $=\textstyle\sum_{i=1}^{n}\phi_{i,k+1}\left(u_{s,k}+\frac{1}{n}(x_{i,k}-y_{i,k})\right),$
$x_{s,k+1}$ $=x_{s,k}+\phi_{s,k+1}u_{s,k},$
where $g_{i}(x_{s,k})$ is a stochastic approximation of the (sub)gradient of $f_{i}(\cdot)$ at $x_{s,k}$.
We note that by construction, since each iteration count $k\in\mathbb{N}$ corresponds to a single event from the set $\{E_{1},...,E_{n},E_{s}\}$, only a single variable in the set $\{\phi_{1,k},...,\phi_{n,k},\phi_{s,k}\}$ is equal to $1$ and the rest are 00. To further simplify the above iterates, let us define the variables $z_{i,k}\triangleq x_{s,k}+u_{s,k}-y_{i,k}$ for $i\in[n]$, and recall that $u_{s,0}=0$ and $y_{i,0}=x_{s,0}$ for $i\in[n]$. We then have $\bar{z}_{k+1}=x_{s,k+1}+u_{k+1}-\bar{y}_{k+1}=x_{s,k}+\left(\phi_{s,k+1}+\sum_{i=1}^{n}\phi_{i,k+1}\right)u_{s,k}-\bar{y}_{k}=\bar{z}_{k}$. Hence, $\bar{z}_{k}=\bar{z}_{0}=0$ for all $k\in\mathbb{N}$, and we can eliminate the aggregator variable by setting $u_{s,k}=\bar{y}_{k}-x_{s,k}$. The final form of the iterates of AREA then becomes
$x_{i,k+1}$ $=x_{i,k}+\phi_{i,k+1}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\right),\quad\forall i\in[n],$ (2)
$y_{i,k+1}$ $=y_{i,k}+\phi_{i,k+1}(x_{i,k}-y_{i,k}),\quad\forall i\in[n],$ (3)
$x_{s,k+1}$ $=x_{s,k}+\phi_{s,k+1}(\bar{y}_{k}-x_{s,k}).$ (4)
We are now ready to list our main theoretical results on the convergence of AREA.
4.1 Main results
We provide theoretical results on the convergence of AREA for two instances of Problem (1): (i) smooth and strongly convex $f_{i}(\cdot)$ for $i\in[n]$ over $\mathcal{X}=\mathbb{R}^{p}$ and (ii) convex nonsmooth $f_{i}(\cdot)$ for $i\in[n]$ over convex and compact $\mathcal{X}\subset\mathbb{R}^{p}$. We first list a number of not necessarily overlapping assumptions on the functions $f_{i}(\cdot)$ of Problem (1).
Assumption 5.
(Smoothness) The function $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ in Problem (1) has $L_{i}$-Lipschitz gradients, namely, $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L_{i}\|x-y\|,$ for all $x,y\in\mathbb{R}^{p}$ and $i\in[n]$.
Assumption 6.
(Strong convexity) The function $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ in Problem (1) is $\mu_{i}$-strongly convex, namely, $f_{i}(y)\geq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{\mu_{i}}{2}\|y-x\|^{2},$ for all $x,y\in\mathbb{R}^{p}$ and $i\in[n]$.
Assumption 7.
(Bounded subgradients) The function $f_{i}:\mathbb{R}^{p}\rightarrow\mathbb{R}$ of Problem (1) is convex for $i\in[n]$ but not necessarily smooth; moreover, there exists a positive constant $C$ such that $\|d_{i}\|\leq C$ for all $d_{i}\in\partial f_{i}(x)$, namely, for all $d_{i}$ such that $f_{i}(y)\geq f_{i}(x)+\langle d_{i},y-x\rangle$ for all $x,y\in\mathbb{R}^{p}$.
Next, we state our main result on the convergence of AREA for strongly convex, smooth functions.
Theorem 8 (Convergence for strongly convex & smooth functions).
Under Assumptions 5 and 6, suppose that ($i$) $\mathcal{X}=\mathbb{R}^{p}$ in Problem (1) and ($ii$) $\alpha_{k}=\alpha\in\left(0,\min_{i}\left\{\frac{2}{\mu_{i}+L_{i}}\right\}\right)$ in Eq. (2) for all $k\in\mathbb{N}_{0}$. Moreover, define the points $x_{i}^{\star}\triangleq x^{\star}-\alpha\nabla f_{i}(x^{\star})\in\mathbb{R}^{p}$ for $i\in[n]$, where $x^{\star}$ is the optimal solution of Problem (1), and the vector $U_{k}\in\mathbb{R}^{2n+1}$ for $k\in\mathbb{N}_{0}$ where $[U_{k}]_{i}\triangleq\mathbb{E}[\|x_{i,k}-x_{i}^{\star}\|^{2}]$ and $[U_{k}]_{n+i}\triangleq\mathbb{E}[\|y_{i,k}-x_{i}^{\star}\|^{2}]$ for $i\in[n]$, and $[U_{k}]_{2n+1}\triangleq\mathbb{E}[\|x_{s,k}-x^{\star}\|^{2}]$. Then for $k\in\mathbb{N}_{0}$, the sequence $\{U_{k}\}$ satisfies the relation
$U_{k}\leq A^{k}U_{0}+\alpha^{2}\sigma^{2}\sum_{t=0}^{k-1}A^{t}q,$ (5)
where $A=\begin{bmatrix}I-P&0&P(1_{n}-\alpha\gamma)\\ P&I-P&0\\ 0&\frac{p_{s}}{n}1_{n}^{\prime}&1-p_{s}\end{bmatrix}\in\mathbb{R}^{(2n+1)\times(2n+1)},\quad q=\begin{bmatrix}P1_{n}\\ 0\\ 0\end{bmatrix}\in\mathbb{R}^{2n+1}$, $P=diag(p_{1},...,p_{n})\in\mathbb{R}^{n\times n}$, $1_{n}$ is the vector of ones of size $n$, $\gamma=[\gamma_{1},...,\gamma_{n}]^{\prime}\in\mathbb{R}^{n}$, $\gamma_{i}=\frac{2\mu_{i}L_{i}}{\mu_{i}+L_{i}}$, and $\sigma^{2}$ is defined in Assumption 1.
Moreover, $\rho(A)<1$, and $\left[\sum_{t=0}^{k-1}A^{t}q\right]_{n+1}\leq\frac{1}{\alpha\bar{\gamma}}$ for $k\in\mathbb{N}_{0}$, where $\bar{\gamma}=\frac{1}{n}\sum_{i=1}^{n}\gamma_{i}$.
The result of Theorem 8 can be interpreted as follows: the expected distance to optimality of the server iterates of AREA at the $k^{th}$ iteration, i.e. $[U_{k}]_{2n+1}$, is characterized by the $(2n+1)^{th}$ rows of the two terms on the right-hand side of (5). The first term, $[A^{k}U_{0}]_{2n+1}$, decays exponentially fast with $k$ with rate $\mathcal{O}(\rho(A))$, while the second term is upper bounded by $\frac{\alpha\sigma^{2}}{\bar{\gamma}}$ for $k\in\mathbb{N}_{0}$. We note that the second term is independent of the probabilities $p_{i}$ for $i\in[n]$ and $p_{s}$, which appear only in the matrix $A$, and hence influence $\rho(A)$. Although it is not possible to find a closed form expression for $\rho(A)$ and obtain tight convergence rates for AREA, when the second term dominates the right-hand side of (5), we obtain error $\mathcal{O}\left(\sigma^{2}/T\right)$ after $T$ iterations with $\alpha=\mathcal{O}(\bar{\gamma}/T)$. This bound is independent of the heterogeneity between the local functions, theoretically validating the effectiveness of our mechanism in eliminating the client drift due to asynchrony.
To the best of our knowledge, convergence guarantees under non-iid data and strong convexity and smoothness are provided only for asynchronous SGD in [8], and the methods AsyncCommSGD [3] and MIFA [17], that employ error-correction schemes similar to ours. The work [8] obtains a convergence rate of $\mathcal{O}\left(\left(n\sigma^{2}+\zeta^{2}\sum_{i=1}^{n}\tau_{i}^{2}\right)/T\right)$, where $\tau_{i}$ is the delay of client $i$ and $\zeta=\max_{i}\|\nabla f_{i}(x)-\nabla f(x)\|$; the last term grows with the heterogeneity between the local datasets and has a quadratic dependency on client delays, suggesting that staleness exacerbates client drift for asynchronous SGD. Conversely, [3, 17] obtain a rate of $\mathcal{O}\left(\sigma^{2}/(nT)+\tau_{\max}^{2}D^{2}/T^{2}\right)$ for the sequence $\{f(\bar{x}_{T})-f^{\star}\}$, where $D$ measures the heterogeneity between the local functions, $\{\bar{x}_{T}\}$ is a sequence of ergodic server iterates, $f^{\star}$ is the optimal value of Problem (1), and $\tau_{\max}$ is the maximum client delay. Unlike [3, 17] (i) we provide a stronger convergence result for the actual, non-ergodic sequence of server iterates, and (ii) completely eliminate the dependency on function heterogeneity and the maximum delay. Although we do not obtain a scaling of $\sigma^{2}$ with $1/n$ like in [3, 17], this is because we consider at most a single local stochastic gradient step at each iteration, unlike [3, 17] that consider only aggregated gradient steps at the server as iterations. Finally, we note that when the true gradients are available ($\sigma=0$ in Assumption 1), AREA converges to the optimal solution of Problem (1) exponentially fast with the number of iterations $k$ (for empirical results on a toy quadratic problem, see Appendix A.2.1).
Remark 9.
In the setting of Example 4, if the client update frequencies $\lambda_{i}$ for $i\in[n]$ are known, one can numerically solve for the aggregation frequency $\lambda_{s}$ which maximizes the convergence rate (i.e., minimizes $\rho(A)$). Moreover, the average time between two consecutive iterations of AREA becomes $1/\bar{\lambda}$, which can be significantly shorter than the time $1/\min_{i}\lambda_{i}$ elapsed between two iterations of synchronous SGD (i.e., waiting for the straggler at each iteration).
We conclude this section with a result on the convergence of AREA when the functions $f_{i}(\cdot)$ of Problem (1) are convex and nonsmooth. In this version of AREA, clients have access only to stochastic approximations of the subgradients of their local functions. Hence, even in the absence of stochastic noise ($\sigma=0$ in Assumption 1), a diminishing stepsize sequence is required to achieve convergence to the set of optimal solutions of Problem (1).
Theorem 10 (Convergence for convex & nonsmooth functions).
Under Assumption 7, suppose that ($i$) the constraint set $\mathcal{X}\subset\mathbb{R}^{p}$ in Problem (1) is convex and compact and ($ii$) the stepsize sequence $\{\alpha_{k}\}$ in Eq. (2) satisfies the following: ($a$) $\alpha_{k+1}\leq\alpha_{k}$ for all $k\in\mathbb{N}_{0}$, ($b$) $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, and ($c$) $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$. Moreover, let $x^{\star}$ be an arbitrary solution of Problem (1), and define $f^{\star}\triangleq f(x^{\star})$ and $V_{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|x_{i,k}-x^{\star}\|^{2}+\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k}-x^{\star}\|^{2}+p_{s}^{-1}\|x_{s,k}-x^{\star}\|^{2}$ for $k\in\mathbb{N}_{0}$. Then the following statement holds for $k=0,1,2,...,$
$\mathbb{E}[V_{k+1}|\mathcal{F}_{k}]\leq V_{k}+\alpha^{2}_{k}\hat{\sigma}^{2}-\frac{2\alpha_{k}}{n}(f(x_{s,k})-f^{\star}),$ (6)
where $\hat{\sigma}^{2}=C^{2}+\sigma^{2}$ and $C^{2}$ and $\sigma^{2}$ are defined in Assumptions 7 and 1, respectively.
In addition, the sequence of server iterates $\{x_{k}\}$ of AREA converges to a solution of Problem (1) with probability $1$.
Theorem 10 establishes the almost sure convergence of the iterates of AREA to an optimal solution of Problem (1), provided the stepsize sequence $\{\alpha_{k}\}$ is square-summable but not summable. Moreover, for general positive sequence $\{\alpha_{k}\}$ we recover the convergence rate of the centralized stochastic projected subgradient method [7]. For a constant stepsize sequence $\alpha_{k}=\alpha>0$ for $k\in\mathbb{N}_{0}$, taking the total expectation on both sides of Eq. (6) and applying recursively yields $\mathbb{E}\left[f\left(\bar{x}_{s,T}\right)-f^{\star}\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(x_{s,k})-f^{\star}]\leq\frac{n}{2}\left(\frac{V_{0}}{\alpha T}+\alpha\hat{\sigma}^{2}\right),$ where $\bar{x}_{s,T}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}x_{s,k}$ and the first inequality holds by Jensen’s inequality. Setting $\alpha=\sqrt{\frac{V_{0}}{\hat{\sigma}^{2}T}}$ and $D_{0}\triangleq\|x_{s,0}-x^{\star}\|$ yields
$\mathbb{E}\left[f\left(\bar{x}_{s,T}\right)-f^{\star}\right]\leq n\hat{\sigma}\sqrt{\frac{V_{0}}{T}}=n\hat{\sigma}D_{0}\sqrt{\frac{p_{s}^{-1}+\frac{2}{n}\sum_{i=1}^{n}p_{i}^{-1}}{T}}.$ (7)
The preceding rate depends only on the average of the inverse probabilities $p_{i}^{-1}$ for $i\in[n]$, rendering AREA less sensitive to excessively fast or slow clients.
To the best of our knowledge, the only known convergence rate for asynchronous SGD under non-iid data and convex, non-smooth functions is $\mathcal{O}\left(\hat{\sigma}D_{0}\sqrt{np_{\max}}/\sqrt{T\sqrt{\bar{p}}}\right)$, where $p_{\max}=\max_{i}p_{i}$ and $\bar{p}=\frac{1}{n}\sum_{i=1}^{n}p_{i}$, as obtained in [14]. The authors of [14] achieve a better dependency on $n$ due to considering each aggregation as an iteration count (we consider both local subgradient steps and aggregations as iterations). However, their rate grows with $p_{\max}$, which may be significantly larger than $\bar{p}$. Moreover, to handle non-iid data, [14] requires a biased sampling scheme where clients are selected according to their importance in the global loss function. In contrast, we allow clients to update the server at their own pace. Finally, in the setting of Example 4, it is possible to obtain a closed form solution for the aggregation frequency $\lambda_{s}$ minimizing $V_{0}$.
5 Numerical Results
In this section we evaluate the empirical performance of AREA. For our simulations we implemented a practical variant of AREA where the aggregation criterion is satisfied if $\Delta\in\mathbb{N}$ client updates have been received (see the discussion in Section 4 on how our convergence analysis can be interpreted as an upper bound for this aggregation policy). The elapsed time between two consecutive updates of client $i\in[n]$ was modeled as an exponential random variable $\text{exp}(\lambda_{i})$. We opted to compare AREA with the following methods: (i) synchronous FedAvg (S-FedAvg) [29], which we use as a synchronous benchmark, (ii) asynchronous FedAvg (AS-FedAvg) (e.g., [20]) and (iii) FedBuff [35], two asynchronous methods sensitive to client drift, and (iv) MIFA [17], an asynchronous method robust to client drift. We note that in every figure in this section, the originally noisy lines have been smoothed to enhance visual clarity while maintaining the underlying data trends.
5.1 Multinomial logistic regression
Let $S_{i}$ be the set of samples locally available to client $i\in[n]$. The local loss function at client $i$ can be expressed as
$\hat{f}_{i}(w^{1},\ldots,w^{C})=\frac{1}{s_{i}}\sum_{j\in S_{i}}\sum_{l=1}^{C}h(j,l)+\frac{\nu}{2}\sum_{l=1}^{C}\|w^{l}\|^{2},$ (8)
where $s_{i}=|S_{i}|$, $h(j,l)=\mathbbm{1}(y_{j}=l)\log\left(\frac{e^{\langle w^{l},x_{j}\rangle}}{\sum_{h=1}^{C}e^{\langle w^{h},x_{j}\rangle}}\right)$, $\mathbbm{1}(\cdot)$ is the indicator function, $(x_{j},y_{j})$ denote the input feature vector $x_{j}\in\mathbb{R}^{d}$ and the output label $y_{j}\in[C]$ of the $j^{th}$ data-point, $w^{l}\in\mathbb{R}^{d}$ for $l\in[C]$ are the model parameters associated with the $l^{th}$ label, and $\nu>0$ is a regularization parameter.
The corresponding optimization can be formulated as $\min_{w^{1},\ldots,w^{C}\in\mathbb{R}^{d}}\quad f(w^{1},\ldots,w^{C})=\sum_{i=1}^{n}\frac{s_{i}}{S}\hat{f}_{i}(w^{1},\ldots,w^{C}),$ where $S=\sum_{i=1}^{n}s_{i}$. We report results for the MNIST dataset [12] and network size $n=128$. We set $\Delta=4$ for all methods, and used samples from a Dirichlet distribution with parameter $a=0.1$ to simulate non-iid local datasets. We tested two different cases of client update frequencies: ($i$) uniform updates, where we set $\lambda_{i}=10$ for all $i\in[n]$, and ($ii$) non-uniform updates, where $\lambda_{i}$ for $i\in[n]$ were drawn from $\mathcal{N}(10,5)$. To evaluate performance, we tracked two metrics over (continuous) time: ($i$) the loss function value at the server model using all training data and ($ii$) the test accuracy of the server model using all test data. The batch size for the local stochastic gradients was set to $b=32$, and the regularization parameter in Problem (8) was set to $\nu=10^{-3}$. For all the methods, we tuned the stepsize $\alpha$ using grid search in the set $\{10^{-3},10^{-2},10^{-1},10^{0},10^{1},10^{2},10^{3},10^{4}\}$ and selected the values resulting in the fastest convergence. We repeated each experiment configuration $10$ times on a CPU cluster with $36$ compute nodes of memory 128GB DDR4 ($\sim 4$ hours per configuration, $\sim 4$ days total computation for the experiments in this section) and extracted the minimum, maximum, and mean values of both selected metrics for the plots. Additional results for non-iid data and $n=4$ can be found in Appendix A.2.2, and our results for iid data can be found in Appendix A.3.
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
Figure 1: Results for $n=128$ clients and one local SGD step per client. Loss function value and test accuracy are plotted for uniform (left) and non-uniform client update frequencies (right) over time. The stepsize values for all methods are listed in the legend.
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
Figure 2: Results for $n=128$ clients and $50$ local SGD steps per client. Loss function value and test accuracy are plotted for uniform (left) and non-uniform client update frequencies (right) over time. The stepsize values for all methods are listed in the legend.
Our results for the execution of a single local SGD step at each client are shown in Fig. 1. The main takeaway from Fig. 1 is that AREA is able to tolerate larger stepsizes and achieve faster and more accurate convergence compared with the other methods for both uniform and non-uniform client update frequencies [^3] We note that the choice of stepsize for FedBuff ($\alpha=10^{1}$) resulted in instabilities at the beginning of the optimization, as seen in Figs. 1(a) and 1(c); however, the next smallest option ($\alpha=10^{0}$) results in significantly slower convergence.. Our results also indicate that MIFA requires small stepsizes to guarantee numerical stability. An additional observation is that the performance of S-FedAvg degrades substantially when the client update frequencies are non-uniform. Finally, although providing theoretical guarantees for AREA in the case where multiple local SGD steps are performed at each client is out of the scope of this work, we tested AREA and the remaining methods with $50$ local SGD steps at each client (Fig. 2). We note that performing multiple SGD steps alleviates communication load but introduces client drift to AREA. In Fig. 2 we observe that as in the single local SGD step case, AREA is able to tolerate larger stepsizes relative to the other methods and achieves faster convergence with respect to testing accuracy (Figs.  2(b) and 2(d)). However, the large stepsize and the introduction of client drift also result in a large error neighborhood in the loss function plots (Figs. 2(a) and 2(c)).
6 Limitations
A significant limitation of our work is that we do not provide theoretical or numerical results for nonconvex functions (e.g. neural networks), which is a more practical setup for FL applications. The study of the convergence properties of AREA in the nonconvex setting is left for future work. Moreover, due to the probabilistic nature of our analysis, which allows for arbitrarily long delays between iterations, it is not possible to readily obtain time complexities unless additional assumptions are adopted on the distributions of client delays, e.g., Example 4.
7 Conclusion
We proposed and analyzed AREA, a new asynchronous stochastic (sub)gradient method that is robust to the client drift caused by heterogeneous client update frequencies and converges under arbitrary delays. For smooth and strongly convex functions, we established the asymptotic convergence of AREA to an error neighborhood whose size depends only on the variance of the stochastic gradient approximations used and is independent of the maximum client delay and the heterogeneity between the local datasets. For constrained problems with convex and non-smooth functions, we recovered the same convergence rate as the centralized projected subgradient method. Our numerical results on a multinomial logistic regression problem suggest that AREA performs favorably compared with state-of-the-art methods, particularly when the local datasets are highly non-iid.
Acknowledgements
This work was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under Contract DE-AC02-06CH11357. We gratefully acknowledge the computing resources provided on Bebop, a high-performance computing cluster operated by the Laboratory Computing Resource Center at Argonne National Laboratory.
Appendix A Appendix
A.1 Proofs and additional theoretical results
A.1.1 Preliminaries
In this section we introduce several definitions and preliminary results that will play a key role in our main analysis of AREA. We begin by listing some relevant definitions.
Definition 11.
(Graph of a matrix)[Definition 6.2.10, [18]] The directed graph of $A\in\mathbb{R}^{n\times n}$, denoted by $G(A)$, is the directed graph on $n$ nodes $P_{1},P_{2},...,P_{n}$ such that there is a directed arc in $G(A)$ from $P_{i}$ to $P_{j}$ if and only if $a_{ij}\neq 0$.
Definition 12.
(Strongly connected graph)[Definition 6.2.13, [18]] A directed graph $G$ is strongly connected if between each pair of distinct nodes $P_{i},P_{j}$ in $\Gamma$ there is a directed path of finite length that begins at $P_{i}$ and ends at $P_{j}$.
Next, we list a number of preliminary lemmas and theorems necessary for our main analysis. The following lemma has been adapted from [Theorem 2.1.14, [33]] and implies that the standard gradient method can be viewed as a contractive mapping.
Lemma 13.
(Contraction) Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a $\mu$-strongly convex function with $L$-Lipschitz gradients. Then for any $x,y\in\mathbb{R}^{n}$ and $\alpha<\frac{2}{\mu+L}$ the following relation holds:
$\begin{split}\|x-\alpha\nabla f(x)-(y-\alpha\nabla f(y))\|^{2}&\leq(1-\alpha\gamma)\|x-y\|^{2},\end{split}$
where $\gamma=\frac{2\mu L}{\mu+L}$.
The next lemma is fundamental for proving the almost sure convergence of sequences of random variables.
Lemma 14.
(Robbins–Siegmund)[Lemma 11, Ch. 2, [38]] Let $\nu_{k},u_{k},\alpha_{k},\beta_{k}$ be nonnegative random variables, and let
$\mathbb{E}[\nu_{k+1}|\mathcal{F}_{k}]\leq(1+\alpha_{k})\nu_{k}-u_{k}+\beta_{k}\quad a.s.$
$\sum_{k=0}^{\infty}\alpha_{k}<\infty\quad a.s.,\quad\sum_{k=0}^{\infty}\beta_{k}<\infty\quad a.s.,$
where $\mathbb{E}[\nu_{k+1}|\mathcal{F}_{k}]$ denotes the conditional mathematical expectation for given $\nu_{0},...,\nu_{k}$, $u_{0},...,u_{k}$, $\alpha_{0},...,\alpha_{k}$, $\beta_{0},...,\beta_{k}$. Then
$\nu_{k}\rightarrow\nu\quad a.s.,\quad\sum_{k=0}^{\infty}u_{k}<\infty\quad a.s.,$
where $\nu\geq 0$ is some random variable.
The following two lemmas have been adapted from [18] and will be used to estimate the convergence rate of AREA under Assumption 6.
Lemma 15.
(Spectral radius)[Theorem 8.1.26, [18]] Let $A\in\mathbb{R}^{n\times n}$ be a nonnegative matrix. Then for any positive vector $v\in\mathbb{R}^{n}$ we have
$\rho(A)\leq\max_{1\leq i\leq n}\frac{1}{v_{i}}\sum_{j=1}^{n}a_{ij}v_{j},$
where $\rho(A)$ is the spectral radius of $A$.
Lemma 16.
(Matrix power)[Corollary 5.6.13, [18]] Let $A\in\mathbb{R}^{n\times n}$ and $\epsilon>0$ be given. There is a constant $C=C(A,\epsilon)$ such that $|(A^{k})_{ij}|\leq C(\rho(A)+\epsilon)^{k}$ for all $k=1,2,...0$ and all $i,j=1,...,n$.
Next, we derive an upper bound for the distance between two points in a compact and convex set after a step of the projected subgradient descent method in the next lemma.
Lemma 17.
(Projected stochastic subgradient descent) Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a convex and compact set and $f:\mathcal{X}\rightarrow\mathbb{R}$ a convex function. Then for any $x,y\in\mathcal{X}$, the following inequality holds:
$\|P_{\mathcal{X}}[x-\alpha g]-y\|^{2}\leq\|x-y\|^{2}+\alpha^{2}\|d\|^{2}+2\alpha(f(y)-f(x))+\alpha^{2}\|\epsilon\|^{2}+2\alpha\langle x-\alpha d-y,\epsilon\rangle,$
where $\alpha$ is a positive constant, $d$ is a subgradient of $f$ at $x$, $g$ is a stochastic approximation of $d$, and $\epsilon=g_{i}-d_{i}$.
Proof.
Using the non-expansiveness property of the projection operator $P_{\mathcal{X}}[\cdot]$ on a convex closed set [e.g., [5], Prop. 2.2.1, p. 88]l, we obtain
$\begin{split}\|P_{\mathcal{X}}[x-\alpha g]-y\|^{2}&\leq\|x-\alpha g-y\|^{2}\\ &=\|x-\alpha d-y\|^{2}+\alpha^{2}\|\epsilon\|^{2}-2\alpha\langle x-\alpha d-y,\epsilon\rangle\\ &=\|x-y\|^{2}+\alpha^{2}\|g\|^{2}+2\alpha\langle d,y-x\rangle+\alpha^{2}\|\epsilon\|^{2}-2\alpha\langle x-\alpha d-y,\epsilon\rangle.\end{split}$
Applying the definition of the subgradient completes the proof. ∎
The next two lemmas are adapted from [18] and reproduced here without proofs; we will use both of these lemmas to derive an extension of Lemma 14 (Robbins–Siegmund) for a system of linear inequalities.
Lemma 18.
(Irreducible matrix)[Theorem 6.2.24, [18]] Let $A\in\mathbb{R}^{n\times n}$. Then A is irreducible if and only if $G(A)$ is strongly connected.
Lemma 19.
(Perron–Frobenius)[Theorem 8.4.4, [18]] Let $A\in\mathbb{R}^{n\times n}$ be irreducible and nonnegative, and suppose that $n\geq 2$. Then ($i$) $\rho(A)>0$, ($ii)$ $\rho(A)$ is an algebraically simple eigenvalue of $A$, ($iii$) there is a unique real vector $x=[x_{i}]$ such that $Ax=\rho(A)x$ and $\sum_{i=1}^{n}x_{i}=1$; this vector is positive, and ($iv$) there is a unique real vector $y=[y_{i}]$ such that $y^{\prime}A=\rho(A)y^{\prime}$ and $\sum_{i=1}^{n}x_{i}y_{i}=1$; this vector is positive.
We are now ready to extend Lemma 14 (Robbins–Siegmund) to accommodate multiple sequences of random variables whose dependencies can be expressed as a linear system of inequalities. Specifically, we aim to form a weighted average of the inequalities in the system and obtain a single inequality of the form required by Lemma 14; Lemma 19 will allow us to define appropriate positive weights for this operation.
Lemma 20.
(Robbins–Siegmund for system of inequalities) Let $a_{i,k}$, $b_{i,k}$, $c_{i,k}$ and $d_{i}(x_{s,k})$ for $i\in[m]$ be nonnegative random variables, and define $A_{k}\triangleq[a_{1,k},...,a_{m,k}]^{\prime}$, $B_{k}\triangleq[b_{1,k},...,b_{m,k}]^{\prime}$, $C_{k}\triangleq[c_{1,k},...,c_{m,k}]^{\prime}$ and $D_{k}\triangleq[d_{1,k},...,d_{m,k}]^{\prime}$. Moreover, suppose that $\sum_{k=0}^{\infty}c_{i,k}<\infty$ for all $i=1,...,m$ and that there exists a nonnegative irreducible matrix $M\in\mathbb{R}^{m\times m}$ and a nonnegative diagonal matrix $\Gamma=diag(\gamma_{i})\in\mathbb{R}^{m\times m}$ such that the following system of inequalities holds for all $k\in\mathbb{N}_{0}$:
$\mathbb{E}[A_{k+1}+B_{k+1}|\mathcal{F}_{k}]\leq MA_{k}+B_{k}+C_{k}-\Gamma D_{k}\quad a.s.$
Then if $\rho(M)=1$, the following statements are true: ($i$) there exist strictly positive weights $w_{i}>0$ for $i\in[m]$ and a random variable $\omega\geq 0$ such that $\sum_{i=1}^{m}w_{i}(a_{i,k}+b_{i,k})\rightarrow\omega$ a.s. and ($ii$) $\sum_{k=0}^{\infty}d_{i}(x_{s,k})<\infty$ a.s. for $i\in\mathcal{I}=\{i\in[m]:\gamma_{i}>0\}$.
Proof.
By the Perron–Frobenius theorem (Lemma 19), $M$ has a strictly positive left eigenvector $w\in\mathbb{R}^{m}$ such that $w^{\prime}M=\rho(M)w^{\prime}=w^{\prime}$. Using the linearity of expectation, we obtain
$\begin{split}\mathbb{E}[w^{\prime}A_{k+1}+w^{\prime}B_{k+1}|\mathcal{F}_{k}]&\leq w^{\prime}MA_{k}+w^{\prime}B_{k}+w^{\prime}C_{k}-w^{\prime}\Gamma D_{k}\\ &=\rho(M)w^{\prime}A_{k}+w^{\prime}B_{k}+w^{\prime}C_{k}-w^{\prime}\Gamma D_{k}\\ &=w^{\prime}A_{k}+w^{\prime}B_{k}+w^{\prime}C_{k}-w^{\prime}\Gamma D_{k}\quad a.s.\end{split}$
We can rewrite the preceding inequality as
$\begin{split}\mathbb{E}\left[\sum_{i=1}^{m}w_{i}(a_{i,k+1}+b_{i,k+1})|\mathcal{F}_{k}\right]&\leq\sum_{i=1}^{m}w_{i}(a_{i,k}+b_{i,k})+\sum_{i=1}^{m}w_{i}c_{i,k}-\sum_{i=1}^{m}w_{i}\gamma_{i}d_{i}(x_{s,k})\quad a.s.\end{split}$
Applying Lemma 14 to the relation above completes the proof. ∎
We have now concluded the preliminary section of our convergence analysis. In the next section we present our main results on the convergence of AREA for strongly convex problems.
A.1.2 Main results for strongly convex & smooth functions
In this section we derive the convergence properties of AREA under Assumptions 5 and 6. We assume that $\alpha_{k}=\alpha>0$ in Eq. (2) for all $k\in\mathbb{N}_{0}$ and that $\mathcal{X}=\mathbb{R}^{p}$ in Problem (1). By Assumption 5 we have $\partial f_{i}(x)=\{\nabla f_{i}(x)\}$ for all $x\in\mathbb{R}^{p}$ and $i\in[n]$. Moreover, under Assumptions 5 and 6, Problem (1) has a unique solution $x^{\star}=\arg\min_{x\in\mathbb{R}^{p}}f(x)$. Let us define the points $x_{i}^{\star}\in\mathbb{R}^{p}$ for $i\in[n]$:
$x_{i}^{\star}\triangleq x^{\star}-\alpha\nabla f_{i}(x^{\star}).$ (9)
By the optimality of $x^{\star}$, we have $\frac{1}{n}\sum_{i=1}^{n}x_{i}^{\star}=x^{\star}$. We will analyze the convergence of AREA in the smooth and strongly convex setting by tracking the progress of a custom Lyapunov function $\{V_{k}\}_{k\in\mathbb{N}_{0}}$, which measures the distance of the global and local iterates of AREA to optimality. Specifically, we construct $V_{k}$ as follows:
$\begin{split}V_{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|x_{i,k}-x_{i}^{\star}\|^{2}+\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k}-x_{i}^{\star}\|^{2}+p_{s}^{-1}\|x_{s,k}-x^{\star}\|^{2}.\end{split}$ (10)
We begin our analysis by establishing an intermediate result on the distance of the local client iterates to optimality in the next lemma.
Lemma 21.
(Client distance to optimality) Suppose that the stepsize in Eq. (2) satisfies $\alpha<\min_{i}\left\{\frac{2}{\mu_{i}+L_{i}}\right\}$. Then the sequences of local iterates $\{x_{i,k}\}$ and $\{y_{i,k}\}$ generated by (2) and (3), respectively, satisfy the following relations for all clients $i\in[n]$ and $k=0,1,2,...$:
$\begin{split}\mathbb{E}[\|x_{i,k+1}-x_{i}^{\star}\|^{2}|\mathcal{F}_{k}]&\leq(1-p_{i})\|x_{i,k}-x_{i}^{\star}\|^{2}+p_{i}(1-\alpha\gamma_{i})\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}p_{i}\sigma^{2},\\ \end{split}$ (11a)
$\mathbb{E}[\|y_{i,k+1}-x_{i}^{\star}\|^{2}|\mathcal{F}_{k}]=(1-p_{i})\|y_{i,k}-x_{i}^{\star}\|^{2}+p_{i}\|x_{i,k}-x_{i}^{\star}\|^{2},$ (11b)
where $p_{i}$ is the probability that the server receives a request by client $i$, the points $x_{i}^{\star}\in\mathbb{R}^{p}$ are defined in (9), $x^{\star}=\arg\min_{x\in\mathbb{R}^{p}}f(x)$, $\gamma_{i}=\frac{2\mu_{i}L_{i}}{\mu_{i}+L_{i}}$ for $i\in[n]$, and $\sigma\geq 0$ is defined in Assumption 1.
Proof.
Subtracting $x_{i}^{\star}$ and taking the squared norm on both sides of (2), we obtain
$\begin{split}\|x_{i,k+1}-x_{i}^{\star}\|^{2}&=(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}\|x_{s,k}-\alpha g_{i}(x_{s,k})-x_{i}^{\star}\|^{2}\\ &=(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}\|x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star}\|^{2}\\ &\quad+\alpha^{2}\phi_{i,k+1}\|g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\|^{2}\\ &\quad+2\alpha\phi_{i,k+1}\langle x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star},g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\rangle\\ &\leq(1-\phi_{i,k+1})\|x_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}(1-\alpha\gamma_{i})\|x_{s,k}-x^{\star}\|^{2}\\ &\quad+\alpha^{2}\phi_{i,k+1}\|g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\|^{2}\\ &\quad+2\alpha\phi_{i,k+1}\langle x_{s,k}-\alpha\nabla f_{i}(x_{s,k})-x_{i}^{\star},g_{i}(x_{s,k})-\nabla f_{i}(x_{s,k})\rangle,\end{split}$
where the first equality follows from the facts that $\phi_{i,k+1}^{2}=\phi_{i,k+1}$ and $(1-\phi_{i,k+1})\phi_{i,k+1}=0$ for all $i\in[n]$ and $k\in\mathbb{N}_{0}$ and the second inequality follows from Lemma 13. Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the preceding relation and applying Assumption 1, we obtain the first claim of this lemma.
To prove the second claim of this lemma, we subtract $x_{i}^{\star}$ from (3) and take the squared norm on both sides to obtain
$\begin{split}\|y_{i,k+1}-x_{i}^{\star}\|^{2}&=(1-\phi_{i,k+1})\|y_{i,k}-x_{i}^{\star}\|^{2}+\phi_{i,k+1}\|x_{i,k}-x_{i}^{\star}\|^{2}.\end{split}$
Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the relation above completes the proof. ∎
In the next lemma we derive an additional intermediate result for the distance between the server iterates of AREA and the optimal solution of Problem (1).
Lemma 22.
(Server distance to optimality) The sequence of server iterates $\{x_{k}\}$ generated by (4) satisfies the following relation for $k=0,1,2...$:
$\mathbb{E}[\|x_{s,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}]\leq(1-p_{s})\|x_{s,k}-x^{\star}\|^{2}+\frac{p_{s}}{n}\sum_{i=1}^{n}\|y_{i,k}-x_{i}^{\star}\|^{2},$
where $p_{s}$ is probability that the aggregation criterion is satisfied.
Proof.
Subtracting $x^{\star}$ from (4) and taking the squared norm on both sides, we obtain
$\begin{split}\|x_{s,k+1}-x^{\star}\|^{2}&=(1-\phi_{s,k+1})\|x_{s,k}-x^{\star}\|^{2}+\phi_{s,k+1}\|\bar{y}_{k}-x^{\star}\|^{2}\\ &=(1-\phi_{s,k+1})\|x_{s,k}-x^{\star}\|^{2}+\phi_{s,k+1}\left\|\frac{1}{n}\sum_{i=1}^{n}(y_{i,k}-x_{i}^{\star})\right\|^{2}\\ &\leq(1-\phi_{s,k+1})\|x_{s,k}-x^{\star}\|^{2}+\frac{\phi_{s,k+1}}{n}\sum_{i=1}^{n}\left\|y_{i,k}-x_{i}^{\star}\right\|^{2},\end{split}$
where the first equality follows from the facts that $\phi_{s,k+1}^{2}=\phi_{s,k+1}$ and $(1-\phi_{s,k+1})\phi_{s,k+1}=0$ for all $k\geq 0$, the second equality from the optimality of $x^{\star}$, and the last inequality from Jensen’s inequality for convex functions and the convexity of the squared norm.
Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the preceding relation completes the proof. ∎
In the next theorem we establish the almost sure convergence of the sequence $\{x_{k}\}$ of server iterates to the optimal solution of Problem (1) under the assumption that the true gradients $\nabla f_{i}(\cdot)$ are available for all $i\in[n]$, that is, $\sigma=0$ in Assumption 1. Specifically, we combine the results of Lemmas 21 and 22 to establish that the sequence $\{V_{k}\}$ defined in (10) is non-increasing with probability $1$. The almost sure convergence of the sequence of server iterates $\{x_{s,k}\}$ to the optimal solution $x^{\star}$ of Problem (1) then follows from Lemma 14.
Theorem 23.
[…]
Proof.
Multiplying Eqs. (11a) and (11b) of Lemma 21 with $p_{i}^{-1}$ and summing them together, we have
$\begin{split}\mathbb{E}[p_{i}^{-1}\left(\|x_{i,k+1}-x_{i}^{\star}\|^{2}+\|y_{i,k+1}-x_{i}^{\star}\|^{2}\right)|\mathcal{F}_{k}]&\leq p_{i}^{-1}\left(\|x_{i,k}-x_{i}^{\star}\|^{2}+\|y_{i,k}-x_{i}^{\star}\|^{2}\right)\\ &\quad-\|y_{i,k}-x_{i}^{\star}\|^{2}+(1-\alpha\gamma_{i})\|x_{s,k}-x^{\star}\|^{2}.\end{split}$
Next, we average the previous relation over all clients $i\in[n]$ and add the result of Lemma 22 multiplied by $p_{s}^{-1}$ to obtain the following inequality:
$\mathbb{E}[V_{k+1}|\mathcal{F}_{k}]\leq V_{k}-\alpha\bar{\gamma}\|x_{s,k}-x^{\star}\|^{2},$
where $\bar{\gamma}=\frac{1}{n}\sum_{i=1}^{n}\gamma_{i}$.
Applying Lemma 14 to the preceding relation yields the following: ($i$) there exists random variable $V^{\star}\geq 0$ such that $\{V_{k}\}\rightarrow V^{\star}$ a.s.; and ($ii$) $\sum_{k=0}^{\infty}\|x_{s,k}-x^{\star}\|^{2}<\infty$ with probability $1$. Applying Lemma 14 to Eq. (11a) of Lemma 21 and using ($ii$), we obtain $\sum_{k=0}^{\infty}\|x_{i,k}-x_{i}^{\star}\|^{2}<\infty$ a.s. for $i\in[n]$, which in turn implies $\sum_{k=0}^{\infty}\|y_{i,k}-x_{i}^{\star}\|^{2}<\infty$ a.s. for $i\in[n]$ by application of Lemma 14 to Eq. (11b) of Lemma 21. Thus, we have $\|x_{s,k}-x^{\star}\|^{2}\rightarrow 0$, $\|x_{i,k}-x_{i}^{\star}\|^{2}\rightarrow 0$ and $\|y_{i,k}-x_{i}^{\star}\|^{2}\rightarrow 0$ a.s., implying $\{V_{k}\}\rightarrow 0$ with probability $1$. ∎
In the preceding theorem, we established the almost sure convergence of the iterates of AREA. An important consequence of our analysis is that the range of acceptable stepsize values for AREA depends exclusively on the strong convexity and Lipschitz gradient parameters of the functions $f_{i}(\cdot)$ of Problem (1) and is independent of the probabilities $p_{s}$ and $p_{i}$ for $i\in[n]$.
We conclude our analysis for strongly convex problems with a result on its convergence in the mean square (m.s.) sense. Although this result is weaker than the result of Theorem 23 and also implied by the latter, it allows us to estimate the convergence rate of AREA. Specifically, the server iterates of AREA can be shown to converge to a neighborhood of the optimal solution of Problem (1) in the m.s. sense with linear rate. We establish the dependency of this neighborhood on the variance of the stochastic gradient approximations and the probabilities $p_{i}$ for $i\in[n]$ and $p_{s}$.
Proof of Theorem 8.
We define the vectors $U_{k}\in\mathbb{R}^{2n+1}$ for $k\in\mathbb{N}_{0}$ as follows:
$U_{j,k}$$\triangleq\begin{cases}\mathbb{E}[\|x_{i,k}-x_{i}^{\star}\|^{2}],\quad&j=i,\quad i=1,...,n\\ \mathbb{E}[\|y_{i,k}-x_{i}^{\star}\|^{2}],\quad&j=n+i,\quad i=1,...,n\\ \mathbb{E}[\|x_{s,k}-x^{\star}\|^{2}],\quad&j=2n+1.\end{cases}$
Taking the total expectation on both sides of the results of Lemmas 21 and 22 yields the following recursion:
$U_{k+1}\leq AU_{k}+\alpha^{2}\sigma^{2}q,$ (12)
where $q=[p_{1},...,p_{n},0^{n+1}]^{\prime}\in\mathbb{R}^{2n+1}$.
In order to prove the convergence of the AREA, it suffices to show that $\rho(A)<1$. Consider the positive vector $v=[v_{1},...,v_{2n+1}]^{\prime}$ satisfying the following properties:
$\begin{cases}v_{j}>(1-\alpha\gamma_{j}),&\quad j=1,...,n\\ v_{j}>v_{j-n},&\quad j=n+1,...,2n\\ v_{j}=1,&\quad j=2n+1\end{cases},\quad\bar{v}\triangleq\frac{1}{n}\sum_{j=n+1}^{2n}v_{j}<1.$
By Lemma 15, the spectral radius of $A$ satisfies
[…]
$\frac{1}{v_{n+i}}\bigg{[}(1-p_{i})v_{i+n}+p_{i}v_{i}\bigg{]},\quad i=1,...,n,\quad 1-p_{s}+p_{s}\bar{v}\bigg{\}}.$
After some algebraic manipulation we obtain
[…]
$1-p_{i}(1-v_{i}v_{i+n}^{-1}),\quad i=1,...,n,\quad 1-p_{s}(1-\bar{v})\bigg{\}}.$
We observe that all the elements of this set are strictly smaller than $1$, and hence $\rho(A)<1$. Moreover, applying Eq. (12) recursively over iterations $0,1,2,...,k$ yields
$U_{k}\leq A^{k}U_{0}+\alpha^{2}\sigma^{2}\sum_{j=0}^{k-1}A^{j}q,\quad k\in\mathbb{N}.$
Let $\tilde{\rho}=\rho(A)+\epsilon$ for some $\epsilon\in(0,1-\rho(A))$. By $\rho(A)<1$, we have $\lim_{k\rightarrow\infty}A^{k}U_{0}=0$; and by Lemma 16, the elements of $A^{k}$ are $\mathcal{O}(\tilde{\rho}^{k})$ for all $k\in\mathbb{N}$, which implies that $A^{k}U_{0}$ diminishes with linear rate.
Since $\rho(A)<1$ for each element $j\in[2n+1]$ of $U_{k}$, we have
$\lim\sup_{k\rightarrow\infty}[U_{k}]_{j}\leq\alpha^{2}\sigma^{2}[(I-A)^{-1}q]_{j},$
where it can be verified that $(I-A)^{-1}$ is a block matrix of the form
$\begin{split}&(I-A)^{-1}=\begin{bmatrix}E&F\\ G&H\end{bmatrix}\quad where\quad E=\begin{bmatrix}P^{-1}&0\\ P^{-1}&P^{-1}\end{bmatrix}+\frac{1}{\alpha\bar{\gamma}n}(1_{2}1_{2}^{\prime}\otimes 1_{n}1_{n}^{\prime}P^{-1})\in\mathbb{R}^{2n\times 2n},\\ &F=\frac{1}{\alpha\bar{\gamma}p_{s}}(1_{2}\otimes(1_{n}-\alpha\gamma))\in\mathbb{R}^{2n\times 1},\quad G=\frac{1}{\alpha\bar{\gamma}}(1_{2}^{\prime}\otimes 1_{n}^{\prime}P^{-1})\in\mathbb{R}^{1\times 2n},\quad H=\frac{1}{\alpha\bar{\gamma}p_{s}}\in\mathbb{R},\end{split}$
and the symbol $\otimes$ denotes the Kronecker product operation.
Multiplying the matrix $(I-A)^{-1}$ with the vector $q$ completes the proof. ∎
We have now concluded our analysis of AREA in the strongly convex setting. In the next subsection we present our theoretical results for constrained problems with convex but not necessarily smooth functions.
A.1.3 Main results for convex problems
In this subsection we derive the convergence properties of AREA under Assumption 7. Moreover, for the remainder of this subsection we assume that the constraint set $\mathcal{X}$ of Problem (1) is convex and compact. Similarly to the preceding subsection, we study the progress of a custom Lyapunov function $\{V_{k}\}_{k\in\mathbb{N}_{0}}$. Let $x^{\star}\in\mathcal{X}$ be an arbitrary optimal solution of Problem (1). We define the sequence $\{V_{k}\}$ as
$V_{k}\triangleq\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|x_{i,k}-x^{\star}\|^{2}+\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k}-x^{\star}\|^{2}+p_{s}^{-1}\|x_{s,k}-x^{\star}\|^{2},$ (13)
where $x_{i,k}$ $y_{i,k}$ and $x_{s,k}$ are the iterates of AREA generated by Eqs. (2)–(4).
Our analysis can be divided into two parts. In the first part, we demonstrate that the local client iterates of AREA converge to the global server iterates with probability $1$. To obtain this result, we construct a suitable Lyapunov function (different from $V_{k}$) using Lemma 20. In the second part, we measure the progress of AREA using the Lyapunov function $V_{k}$ of Eq. (13) and derive the almost sure convergence of the method by applying Lemma 14.
We begin our analysis by bounding the difference between the evaluations of the local function $f_{j}(\cdot)$ at the server iterates and the local iterates of client $i$ for all $i,j\in[n]$.
Lemma 24.
(Client-server objective difference) Let $f_{j}^{\star}=\min_{x}f_{j}(x)$ for $j\in[n]$. The following relations hold for all $i\in[n]$ and $k\in\mathbb{N}_{0}$:
$\begin{split}f_{j}(x_{i,k})-f_{j}(x_{s,k})&\leq p_{i}^{-1}(f_{j}(x_{i,k})-f_{j}^{\star})-p_{i}^{-1}\mathbb{E}[f_{j}(x_{i,k+1})-f_{j}^{\star}|\mathcal{F}_{k}]+\alpha_{k}\sigma_{1}^{2},\end{split}$ (14a)
$\begin{split}f_{j}(y_{i,k})-f_{j}(x_{s,k})&\leq p_{i}^{-1}(f_{j}(y_{i,k})+f_{j}(x_{i,k})-2f_{j}^{\star})\\ &\quad-p_{i}^{-1}\mathbb{E}[f_{j}(y_{i,k+1})+f_{j}(x_{i,k+1})-2f_{j}^{\star}|\mathcal{F}_{k}]+\alpha_{k}\sigma_{1}^{2},\end{split}$ (14b)
where $\sigma_{1}^{2}=C(C+\sigma)$, and $C>0$ and $\sigma>0$ are defined in Assumptions 7 and 1, respectively.
Proof.
Evaluating the function $f_{j}(x)$ on $x_{i,k+1}$ of Eq. (2) yields
$\begin{split}f_{j}(x_{i,k+1})&=(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}f_{j}(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})])\\ &\leq(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}\bigg{(}f_{j}(x_{s,k})\\ &\quad-\langle d_{j}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\right),x_{s,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\rangle\bigg{)},\end{split}$
where $d_{j}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\right)$ is a subgradient of $f_{j}(\cdot)$ at $P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]$.
Applying the Cauchy–Schwarz inequality to the inner product term in this relation yields
$\begin{split}f_{j}(x_{i,k+1})&\leq(1-\phi_{i,k+1})f_{j}(x_{i,k})+\phi_{i,k+1}\bigg{(}f_{j}(x_{s,k})\\ &\quad+\left\|d_{j}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\right)\right\|\|x_{s,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]\|\bigg{)}\\ &\leq f_{j}(x_{i,k})+\phi_{i,k+1}(f_{j}(x_{s,k})-f_{j}(x_{i,k}))+\alpha_{k}\phi_{i,k+1}C(C+\|g_{i}(x_{s,k})-d_{i}(x_{s,k})\|),\end{split}$
where $d_{i}(x_{s,k})$ is a subgradient of $f_{i}(\cdot)$ at $x_{s,k}$. The last inequality in the preceding relation follows from the non-expansiveness of the projection operator $P_{\mathcal{X}}[\cdot]$ and Assumption 7.
After subtracting $f_{j}^{\star}$ from both sides of the preceding relation, taking the expectation conditional on $\mathcal{F}_{k}$, and applying Assumption 1 (note that $\mathbb{E}[\|g_{i}(x_{s,k})-d_{i}(x_{s,k})\||\mathcal{F}_{k}]\leq\sigma$ by Jensen’s inequality and the concavity of the square root function), we obtain
$\begin{split}\mathbb{E}[f_{j}(x_{i,k+1})-f_{j}^{\star}|\mathcal{F}_{k}]&\leq f_{j}(x_{i,k})-f_{j}^{\star}+p_{i}(f_{j}(x_{s,k})-f_{j}(x_{i,k}))+\alpha_{k}p_{i}C(C+\sigma).\end{split}$
Multiplying both sides of the relation above with $p_{i}^{-1}$, rearranging the terms, and applying the definition of $\sigma_{1}$, we have the first result of this lemma.
Evaluating the function $f_{j}$ on $y_{i,k+1}$ of Eq. (3) yields
$\begin{split}f_{j}(y_{i,k+1})&=(1-\phi_{i,k+1})f_{j}(y_{i,k})+\phi_{i,k+1}f_{j}(x_{i,k})\\ &=f_{j}(y_{i,k})+\phi_{i,k+1}(f_{j}(x_{i,k})-f_{j}(x_{s,k})+f_{j}(x_{s,k})-f_{j}(y_{i,k})),\end{split}$
where we added and subtracted the quantity $\phi_{i,k+1}f_{j}(x_{s,k})$ to get the last equality.
Subtracting the quantity $f_{j}^{\star}$ from both sides of the relation above and taking the expectation conditional on $\mathcal{F}_{k}$, we obtain
$\begin{split}\mathbb{E}[f_{j}(y_{i,k+1})-f_{j}^{\star}|\mathcal{F}_{k}]&=f_{j}(y_{i,k})-f_{j}^{\star}+p_{i}(f_{j}(x_{i,k})-f_{j}(x_{s,k})+f_{j}(x_{s,k})-f_{j}(y_{i,k})).\end{split}$
Multiplying both sides of the relation above with $p_{i}^{-1}$, applying Eq. (14a) and rearranging the terms, we conclude the proof. ∎
In the next lemma we derive upper bounds for the distances of the local iterates of client $i$ and the global server iterates for all $i\in[n]$.
Lemma 25.
(Client-server distance) The distances between the the server iterates $x_{s,k}$, generated by (4), and the local iterates $x_{i,k}$ and $y_{i,k}$ at client $i$, generated by (2) and (3), respectively, satisfy the following inequalities for $i\in[n]$ and $k\in\mathbb{N}_{0}$:
$\begin{split}\mathbb{E}[\|x_{i,k+1}-x_{s,k+1}\|^{2}|\mathcal{F}_{k}]&\leq(1-p_{s})\|x_{i,k}-x_{s,k}\|^{2}+\frac{p_{s}}{n}\sum_{j=1}^{n}\|x_{i,k}-y_{j,k}\|^{2}+\alpha^{2}_{k}p_{i}\hat{\sigma}^{2}-p_{i}D^{x}_{i,k},\end{split}$ (15a)
$\begin{split}\mathbb{E}[\|y_{i,k+1}-x_{s,k+1}\|^{2}|\mathcal{F}_{k}]&\leq\left(1-p_{i}-\frac{n-1}{n}p_{s}\right)\|y_{i,k}-x_{s,k}\|^{2}\\ &\quad+\frac{p_{s}}{n}\sum_{j\neq i}\|y_{i,k}-y_{j,k}\|^{2}+p_{i}\|x_{i,k}-x_{s,k}\|^{2}-\frac{p_{s}}{n}D^{y}_{i,k},\end{split}$ (15b)
where $\hat{\sigma}^{2}=C^{2}+\sigma^{2}$, $C>0$ and $\sigma>0$ are defined in Assumptions 7 and 1, $D^{x}_{i,k}=\|x_{i,k}-x_{s,k}\|^{2}$, and $D^{y}_{i,k}=\|y_{i,k}-x_{s,k}\|^{2}$.
Proof.
Let $\epsilon_{i,k}=g_{i}(x_{s,k})-d_{i}(x_{s,k})$, where $d_{i}(x_{s,k})$ is a subgradient of $f_{i}(\cdot)$ at $x_{s,k}$. Subtracting (4) from (2) and taking the squared norm on both sides, we get
$\begin{split}\|x_{i,k+1}-x_{s,k+1}\|^{2}&=\phi_{s,k+1}\|x_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{s,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2}\\ &\leq\phi_{s,k+1}\|x_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}(\alpha^{2}_{k}C^{2}+\alpha^{2}_{k}\|\epsilon_{i,k}\|^{2}+2\alpha_{k}^{2}\langle d_{i}(x_{s,k}),\epsilon_{i,k}\rangle)\\ &\quad+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2}\\ &\leq\frac{\phi_{s,k+1}}{n}\sum_{j=1}^{n}\|x_{i,k}-y_{j,k}\|^{2}+\phi_{i,k+1}\bigg{(}\alpha^{2}_{k}C^{2}+\alpha^{2}_{k}\|\epsilon_{i,k}\|^{2}\\ &\quad+2\alpha_{k}^{2}\langle d_{i}(x_{s,k}),\epsilon_{i,k}\rangle\bigg{)}+(1-\phi_{i,k+1}-\phi_{s,k+1})\|x_{i,k}-x_{s,k}\|^{2},\end{split}$
where we obtain the first inequality after combining Lemma 17 and Assumption 7 and the last inequality from Jensen’s inequality and the convexity of the squared norm.
Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the preceding relation and applying Assumption 1 and the definition of $D^{x}_{i,k}$, we obtain the first result of this lemma.
Subtracting (4) from (3) and taking the squared norm on both sides, we get
$\begin{split}\|y_{i,k+1}-x_{s,k+1}\|^{2}&=\phi_{s,k+1}\|y_{i,k}-\bar{y}_{k}\|^{2}+\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}\\ &\quad+(1-\phi_{s,k+1}-\phi_{i,k+1})\|y_{i,k}-x_{s,k}\|^{2}\\ &\leq\frac{\phi_{s,k+1}}{n}\sum_{j=1}^{n}\|y_{i,k}-y_{j,k}\|^{2}+\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}\\ &\quad+(1-\phi_{s,k+1}-\phi_{i,k+1})\|y_{i,k}-x_{s,k}\|^{2},\end{split}$
where we obtain the last inequality from Jensen’s inequality and the convexity of the squared norm.
Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the above relation and applying the definition of $D^{y}_{i,k}$, we complete the proof. ∎
In the next lemma we derive upper bounds for the distances between the local iterates of any two clients $i,j\in[n]$ with $i\neq j$ and the distances between the iterates locally maintained by client $i$ for $i\in[n]$.
Lemma 26.
(Interagent distance) Suppose that $\alpha_{k+1}\leq\alpha_{k}$ in Eq. (2) for all $k\in\mathbb{N}_{0}$. For any clients $i,j\in[n]$ with $i\neq j$ and $k\in\mathbb{N}_{0}$, the following hold:
$\begin{split}&\mathbb{E}\bigg{[}\|x_{i,k+1}-y_{j,k+1}\|^{2}+\alpha_{k+1}\left(B^{x}_{ij,k+1}+B^{y}_{ij,k+1}\right)|\mathcal{F}_{k}\bigg{]}\leq p_{i}\|x_{s,k}-y_{j,k}\|^{2}\\ &\quad+p_{j}\|x_{i,k}-x_{j,k}\|^{2}+(1-p_{i}-p_{j})\|x_{i,k}-y_{j,k}\|^{2}+\alpha_{k}^{2}p_{i}\sigma_{3}^{2}+\alpha_{k}\left(B^{x}_{ij,k}+B^{y}_{ij,k}\right),\end{split}$ (16a)
$\begin{split}&\mathbb{E}\bigg{[}\|x_{i,k+1}-x_{j,k+1}\|^{2}+\alpha_{k+1}\left(B^{x}_{ij,k+1}+B^{x}_{ji,k+1}\right)|\mathcal{F}_{k}\bigg{]}\leq p_{i}\|x_{j,k}-x_{s,k}\|^{2}+p_{j}\|x_{i,k}-x_{s,k}\|^{2}\\ &\quad+(1-p_{i}-p_{j})\|x_{i,k}-x_{j,k}\|^{2}+\alpha_{k}^{2}(p_{i}+p_{j})\sigma^{2}_{3}+\alpha_{k}\left(B^{x}_{ij,k}+B^{x}_{ji,k}\right),\end{split}$ (16b)
$\begin{split}\mathbb{E}[\|y_{i,k+1}-y_{j,k+1}\|^{2}|\mathcal{F}_{k}]&=p_{i}\|x_{i,k}-y_{j,k}\|^{2}+p_{j}\|y_{i,k}-x_{j,k}\|^{2}+(1-p_{i}-p_{j})\|y_{i,k}-y_{j,k}\|^{2},\end{split}$ (16c)
where $B^{x}_{ij,\tau}=2p_{i}p_{j}^{-1}(f_{i}(x_{j,\tau})-f_{i}^{\star})$ , $B^{y}_{ij,\tau}=2p_{i}p_{j}^{-1}(f_{i}(y_{j,\tau})-f_{i}^{\star})$ for any $i,j\in[n]$ and $\tau\in\mathbb{N}_{0}$, $\sigma_{3}^{2}=\hat{\sigma}^{2}+2\sigma_{1}^{2}$, $\hat{\sigma}^{2}$ is defined in Lemma 25, and $\sigma_{1}^{2}$ is defined in Lemma 24.
Moreover, the distance between the local iterates at client $i$ for all $i\in[n]$ and $k\in\mathbb{N}_{0}$ satisfies
$\begin{split}\mathbb{E}[\|x_{i,k+1}-y_{i,k+1}\|^{2}+\alpha_{k+1}B^{x}_{ii,k+1}|\mathcal{F}_{k}]&\leq p_{i}\|x_{i,k}-x_{s,k}\|^{2}+(1-p_{i})\|x_{i,k}-y_{i,k}\|^{2}\\ \ &\quad+\alpha_{k}B^{x}_{ii,k+1}+\alpha_{k}^{2}p_{i}\sigma_{3}^{2}.\end{split}$ (17)
Proof.
Let $\epsilon_{i,k}=g_{i}(x_{s,k})-d_{i}(x_{s,k})$, where $d_{i}(x_{s,k})$ is a subgradient of $f_{i}(\cdot)$ at $x_{s,k}$ for $i\in[n]$. Subtracting (3) applied for the $j^{th}$ client from (2) applied for the $i^{th}$ client and taking the squared norm on both sides, we get
$\begin{split}\|x_{i,k+1}-y_{j,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-y_{j,k}\|^{2}+\phi_{j,k+1}\|x_{i,k}-x_{j,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-y_{j,k}\|^{2}\\ &\leq\phi_{i,k+1}\|y_{j,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})\\ &\quad+2\alpha_{k}\phi_{i,k+1}(f_{i}(y_{j,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-y_{j,k},\epsilon_{i,k}\rangle\\ &\quad+\phi_{j,k+1}\|x_{i,k}-x_{j,k}\|^{2}+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-y_{j,k}\|^{2},\end{split}$
where the last inequality follows from Lemma 17 and Assumption 7.
Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the preceding relation and applying Assumption 1 and Eq. (14b) of Lemma 24, we have
$\begin{split}&\mathbb{E}[\|x_{i,k+1}-y_{j,k+1}\|^{2}|\mathcal{F}_{k}]\leq p_{i}\|y_{j,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}p_{i}(C^{2}+\sigma^{2}+2\sigma_{1}^{2})\\ &\quad+p_{j}\|x_{i,k}-x_{j,k}\|^{2}+(1-p_{i}-p_{j})\|x_{i,k}-y_{j,k}\|^{2}\\ &\quad+2\alpha_{k}p_{i}p_{j}^{-1}\left\{f_{i}(y_{j,k})+f_{i}(x_{j,k})-2f_{i}^{\star}-\mathbb{E}[f_{i}(y_{j,k+1})+f_{i}(x_{j,k+1})-2f_{i}^{\star}|\mathcal{F}_{k}]\right\}.\end{split}$
Using the fact that $\alpha_{k+1}\leq\alpha_{k}$ and applying the definitions of $B^{x}_{ij,\tau}$, $B^{y}_{ij,\tau}$ and $\sigma_{3}^{2}$ and rearranging, we obtain the first result of this lemma.
Similarly, subtracting (2) applied for the $j$th client from (2) applied for the $i$th client and taking the squared norm on both sides, we have
$\begin{split}\|x_{i,k+1}-x_{j,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{j,k}\|^{2}\\ &\quad+\phi_{j,k+1}\|x_{i,k}-P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{j}(x_{s,k})]\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k+1}-x_{j,k+1}\|^{2}\\ &\leq\phi_{i,k+1}\|x_{j,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})\\ &\quad+2\alpha_{k}\phi_{i,k+1}(f_{i}(x_{j,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x_{j,k},\epsilon_{i,k}\rangle\\ &\quad+\phi_{j,k+1}\|x_{i,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{j,k+1}(C^{2}+\|\epsilon_{j,k}\|^{2})\\ &\quad+2\alpha_{k}\phi_{j,k+1}(f_{j}(x_{i,k})-f_{j}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{j,k+1}\langle x_{s,k}-\alpha_{k}d_{j}(x_{s,k})-x_{i,k},\epsilon_{j,k}\rangle\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|x_{i,k}-x_{j,k}\|^{2},\end{split}$
where we obtain the last inequality by applying Lemma 17 and Assumption 7. Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the relation above and applying Assumption 1, Lemma 24, the fact that $\alpha_{k+1}\leq\alpha_{k}$, the definitions of $B^{x}_{ij,\tau}$ and $\sigma_{3}^{2}$ and rearranging, we obtain the second result of this lemma.
Subtracting (3) applied for the $j$th client from (3) applied for the $i$th client and taking the squared norm on both sides, we have
$\begin{split}\|y_{i,k+1}-y_{j,k+1}\|^{2}&=\phi_{i,k+1}\|x_{i,k}-y_{j,k}\|^{2}+\phi_{j,k+1}\|y_{i,k}-x_{j,k}\|^{2}\\ &\quad+(1-\phi_{i,k+1}-\phi_{j,k+1})\|y_{i,k}-y_{j,k}\|^{2}.\end{split}$
where we obtain the last inequality by applying Lemma 17 and Assumption 7. Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the relation above, we obtain the third result of this lemma.
Subtracting (3) applied for the $i$th client from (2) applied for the $i$th client and taking the squared norm on both sides, we have
$\begin{split}\|x_{i,k+1}-y_{i,k+1}\|^{2}&=\phi_{i,k+1}\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\|^{2}+(1-\phi_{i,k+1})\|x_{i,k}-y_{i,k}\|^{2}\\ &\leq\phi_{i,k+1}\|x_{i,k}-x_{s,k}\|^{2}+\alpha_{k}^{2}\phi_{i,k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})\\ &\quad+2\alpha_{k}\phi_{i,k+1}(f_{i}(x_{i,k})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\phi_{i,k+1}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x_{i,k},\epsilon_{i,k}\rangle+(1-\phi_{i,k+1})\|x_{i,k}-y_{i,k}\|^{2},\end{split}$
where the last inequality follows from Lemma 17 and Assumption 7. Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the preceding relation, and applying Assumption 1, Lemma 24, $\alpha_{k+1}\leq\alpha_{k}$, and the definitions of $B^{x}_{ii,\tau}$, and $\sigma_{3}^{2}$ and rearranging, we conclude the proof. ∎
We are now ready to combine the results of Lemmas 2426 into a weighted average that will serve as a Lyapunov function measuring the distances between the local and global iterates of AREA. Rather than explicitly defining the weights for each inequality, it is sufficient to prove that they exist and are strictly positive. The almost sure convergence of the local client iterates to the global server iterates then follows from Lemma 20.
Theorem 27.
(Client-server convergence) Suppose that the stepsize sequence $\{\alpha_{k}\}$ in Eq. (2) satisfies the following: ($i$) $\alpha_{k+1}\leq\alpha_{k}$ for all $k\in\mathbb{N}$, ($ii$) $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, and ($iii$) $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$. Then the distances of the local client iterates $x_{i,k}$ and $y_{i,k}$ generated by (2) and (3), respectively, to the server iterates $x_{s,k}$ produced by (4) converge to zero almost surely for all $i\in[n]$, namely,,
$\sum_{k=0}^{\infty}\|x_{i,k}-x_{s,k}\|^{2}<\infty\quad a.s.,\quad\sum_{k=0}^{\infty}\|y_{i,k}-x_{s,k}\|^{2}<\infty\quad a.s.\quad\forall i\in[n].$
Proof.
We define the following quantities for any pair of clients $i,j\in[n]$ and $k\in\mathbb{N}_{0}$:
$A^{i}_{xs}(k):=\|x_{i,k}-x_{s,k}\|^{2},\quad A^{i}_{ys}(k):=\|y_{i,k}-x_{s,k}\|^{2},\quad A^{ij}_{xy}(k):=\|x_{i,k}-y_{j,k}\|^{2},$
$A^{ij}_{xx}(k):=\|x_{i,k}-x_{j,k}\|^{2},\quad A^{ij}_{yy}(k):=\|y_{i,k}-y_{j,k}\|^{2}.$
Moreover, let $A_{k}\in\mathbb{R}^{n+2n^{2}}$ be a vector containing the following elements in arbitrary order:
  • $A^{i}_{xs}(k)$ for all $i\in[n]$,
  • $A^{i}_{ys}(k)$ for all $i\in[n]$,
  • $A^{ij}_{xy}(k)$ for all $(i,j)\in[n]\times[n]$,
  • $A^{ij}_{xx}(k)$ for all $i\in[n]$ and $j\in[n]:j>i$, and
  • $A^{ij}_{yy}(k)$ for all $i\in[n]$ and $j\in[n]:j>i$.
Equations (15a) and (15b) of Lemma 25 and (16a)–(17) of Lemma 26 can be combined to form the following system of inequalities:
$\mathbb{E}[A_{k+1}+\alpha_{k+1}B_{k+1}|\mathcal{F}_{k}]\leq MA_{k}+\alpha_{k}B_{k}+\alpha_{k}^{2}C_{k}-\Gamma D_{k},$ (18)
where the quantities in the preceding relation are defined as follows:
  • $M\in\mathbb{R}^{(n+2n^{2})\times(n+2n^{2})}$ is a matrix containing the nonnegative coefficients of elements of $A_{k}$, that is, the coefficients of the squared normed terms on the right-hand side of Eqs.  (15a) and (15b) of Lemma 25 and (16a)–(17) of Lemma 26.
  • $B_{k}\in\mathbb{R}^{n+2n^{2}}$ ($B_{k+1}$, resp.) contains the terms $B^{x}_{ij,k}$ and $B^{y}_{ij,k}$ ($B^{x}_{ij,k+1}$ and $B^{y}_{ij,k+1}$ resp.) in Eqs.  (16a), (16b), and (17) and zeros in the remaining entries (we note that $B^{x}_{ij,\tau}\geq 0$ and $B^{y}_{ij,\tau}\geq 0$ for all $i,j\in[n]$ and $\tau\geq 0$).
  • $C_{k}\in\mathbb{R}^{n+2n^{2}}$ contains the diminishing terms dependent on $\hat{\sigma}^{2}$ and $\sigma_{3}^{2}$ in Eq. (15a) of Lemma 25 and Eqs.  (16a), (16b), and (17) of Lemma 26 and zeros in the remaining entries.
  • $D_{k}\in\mathbb{R}^{n+2n^{2}}$ contains the terms $D^{x}_{i,k}$ and $D^{y}_{i,k}$ in Eqs. (15a) and (15b), respectively, and zeros in the remaining entries.
  • $\Gamma\in\mathbb{R}^{(n+2n^{2})\times(n+2n^{2})}$ is a nonnegative diagonal matrix containing the positive coefficients of the terms in $D^{x}_{i,k}$ and $D^{y}_{i,k}$ in Eqs. (15a) and (15b) and zero otherwise.
The matrix $M$ is nonnegative and row-stochastic, and therefore $\rho(M)\leq 1$ by Lemma 15. Moreover, there is a directed path between every two nodes of the graph $G(M)$ of $M$, and hence $M$ is irreducible. To make the last argument more obvious, we have plotted the subgraph of $G(M)$ for any two clients $i,j\in[n]$ with $i\neq j$ in Fig. 3, which is obviously strongly connected. Thus, by Lemma 19 we have $\rho(M)=1$. Applying Lemma 20 to (18) completes the proof.
Figure 3: Strongly connected subgraph of $G(M)$ for any $i,j\in[n]$ with $i\neq j$
We have now concluded with the first part of our analysis, where we demonstrated that the local client iterates converge to the global server iterates with probability $1$. In the remainder of this subsection we prove that the local and global iterates of AREA converge to the set of optimal solutions of Problem (1). We first bound the distance of the local client iterates to optimality in the next lemma.
Lemma 28.
(Client distance to optimality) Let $x^{\star}$ be an arbitrary optimal solution of Problem (1), and recall that $\hat{\sigma}^{2}=C^{2}+\sigma^{2}$, where $C^{2}$ and $\sigma^{2}$ are defined in Assumptions 7 and 1, respectively. The sequences of local client iterates $\{x_{i,k}\}$ and $\{y_{i,k}\}$ produced by Eq. (2) and (3), respectively, satisfy the following relations for all $i\in[n]$ and $k\in\mathbb{N}_{0}$:
$\begin{split}\mathbb{E}[\|x_{i,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}]&\leq(1-p_{i})\|x_{i,k}-x^{\star}\|^{2}+p_{i}\|x_{s,k}-x^{\star}\|^{2}\\ &\quad+p_{i}\alpha^{2}_{k}\hat{\sigma}^{2}+2p_{i}\alpha_{k}(f_{i}(x^{\star})-f_{i}(x_{s,k})),\end{split}$ (19)
$\begin{split}\mathbb{E}[\|y_{i,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}]&=(1-p_{i})\|y_{i,k}-x^{\star}\|^{2}+p_{i}\|x_{i,k}-x^{\star}\|^{2}.\end{split}$ (20)
Proof.
Subtracting $x^{\star}$ from Eq. (2) and taking the squared norm on both sides, we have
$\begin{split}\|x_{i,k+1}-x^{\star}\|^{2}&=(1-\phi_{i,k+1})\|x_{i,k}-x^{\star}\|^{2}+\phi_{i,k+1}\left\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x^{\star}\right\|^{2}\\ &=(1-\phi_{i,k+1})\|x_{i,k}-x^{\star}\|^{2}+\phi_{i,k+1}\left\|P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-P_{\mathcal{X}}[x^{\star}]\right\|^{2}\\ &\leq(1-\phi_{i,k+1})\|x_{i,k}-x^{\star}\|^{2}+\phi_{i,k+1}\left\|x_{s,k}-\alpha_{k}g_{i}(x_{s,k})-x^{\star}\right\|^{2},\end{split}$
where we used the fact that $x^{\star}\in\mathcal{X}$ to get the first equality and the non-expansiveness property of the projection operator on a convex closed set.
Let $\epsilon_{i,k}=g_{i}(x_{s,k})-d_{i}(x_{s,k})$, where $d_{i}(x_{s,k})$ is a subgradient of $f_{i}(\cdot)$ at $x_{s,k}$. For the second term in the relation above we have
$\begin{split}\|x_{s,k}-\alpha_{k}g_{i}(x_{s,k})-x^{\star}\|^{2}&=\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}_{k}(\|d_{i}(x_{s,k})\|^{2}+\|\epsilon_{i,k}\|^{2})\\ &\quad+2\alpha_{k}\langle d_{i}(x_{s,k}),x^{\star}-x_{s,k}\rangle-2\alpha_{k}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x^{\star},\epsilon_{i,k}\rangle\\ &\leq\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}_{k+1}(C^{2}+\|\epsilon_{i,k}\|^{2})+2\alpha_{k}(f_{i}(x^{\star})-f_{i}(x_{s,k}))\\ &\quad-2\alpha_{k}\langle x_{s,k}-\alpha_{k}d_{i}(x_{s,k})-x^{\star},\epsilon_{i,k}\rangle,\end{split}$
where we used Assumption 7 and the definition of the subgradient to obtain the second inequality.
Combining the two preceding relations, taking the expectation conditional on $\mathcal{F}_{k}$ and applying Assumption 1, we obtain the first result of this lemma.
To obtain the second result, we subtract $x^{\star}$ from Eq. (3) and take the squared norm on both sides:
$\begin{split}\|y_{i,k+1}-x^{\star}\|^{2}&=(1-\phi_{i,k+1})\|y_{i,k}-x^{\star}\|^{2}+\phi_{i,k+1}\|x_{i,k}-x^{\star}\|^{2}.\end{split}$
Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the relation above concludes the proof. ∎
In the next lemma we derive one more intermediate result on the distance of the server iterates to optimality.
Lemma 29.
(Server distance to optimality) Let $x^{\star}$ be an arbitrary optimal solution of Problem (1). The sequence of server iterates $\{x_{s,k}\}$ produced by Eq. (4) satisfies the following relation for $k\in\mathbb{N}_{0}$
$\begin{split}\mathbb{E}[\|x_{s,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}]&\leq(1-p_{s})\|x_{s,k}-x^{\star}\|^{2}+\frac{p_{s}}{n}\sum_{i=1}^{n}\|y_{i,k}-x^{\star}\|^{2}.\end{split}$
Proof.
Subtracting $x^{\star}$ from Eq. (4) and taking the squared norm on both sides, we have
$\begin{split}\|x_{s,k+1}-x^{\star}\|^{2}&=(1-\phi_{s,k+1})\|x_{s,k}-x^{\star}\|^{2}+\phi_{s,k+1}\|\bar{y}_{k}-x^{\star}\|^{2}\\ &\leq(1-\phi_{s,k+1})\|x_{s,k}-x^{\star}\|^{2}+\phi_{s,k+1}\frac{1}{n}\sum_{i=1}^{n}\|y_{i,k}-x^{\star}\|^{2},\end{split}$
where we used the standard inequality ±$\pm 2\langle a,b\rangle\leq\|a\|^{2}+\|b\|^{2}$ that holds for any two vectors $a,b$ to obtain the last inequality.
Taking the expectation conditional on $\mathcal{F}_{k}$ on both sides of the preceding relation concludes the proof. ∎
We conclude our convergence analysis of AREA under Assumption 7 with our main theorem listed below. Specifically, we combine the results of Lemmas 28 and 29 and apply Lemma 14 to show that the Lyapunov function $V_{k}$ defined in (13) converges with probability $1$; the convergence of the global server iterates (and, by extension, of the local client iterates due to Theorem 27) to the set of optimal solutions of Problem (1) then follows.
Proof of Theorem 10.
Multiplying Eq. (19) of Lemma 28 with $p_{i}^{-1}/n$ and summing over all $i\in[n]$, we have
$\begin{split}\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|x_{i,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}\right]&\leq\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|x_{i,k}-x^{\star}\|^{2}+\|x_{s,k}-x^{\star}\|^{2}+\alpha^{2}_{k}\hat{\sigma}^{2}\\ &\quad+\frac{2\alpha_{k}}{n}(f^{\star}-f(x_{s,k}))-\frac{1}{n}\sum_{i=1}^{n}\|x_{i,k}-x^{\star}\|^{2}.\end{split}$
Similarly, multiplying Eq. (20) of Lemma 28 with $p_{i}^{-1}/n$ and summing over $i=1,...,n$, we have
$\begin{split}\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}\right]&\leq\frac{1}{n}\sum_{i=1}^{n}p_{i}^{-1}\|y_{i,k}-x^{\star}\|^{2}+\frac{1}{n}\sum_{i=1}^{n}\|x_{i,k}-x^{\star}\|^{2}\\ &\quad-\frac{1}{n}\sum_{i=1}^{n}\|y_{i,k}-x^{\star}\|^{2}.\end{split}$
Multiplying the result of Lemma 29 with $p_{s}^{-1}$ yields
$\begin{split}\mathbb{E}[p_{s}^{-1}\|x_{s,k+1}-x^{\star}\|^{2}|\mathcal{F}_{k}]&\leq p_{s}^{-1}\|x_{s,k}-x^{\star}\|^{2}+\frac{1}{n}\sum_{i=1}^{n}\|y_{i,k}-x^{\star}\|^{2}-\|x_{s,k}-x^{\star}\|^{2}.\end{split}$
After adding together the 3 preceding relations we obtain
$\begin{split}\mathbb{E}[V_{k+1}|\mathcal{F}_{k}]&\leq V_{k}+\alpha^{2}_{k}\hat{\sigma}^{2}-\frac{2\alpha_{k}}{n}(f(x_{s,k})-f^{\star}).\end{split}$ (21)
Applying Lemma 14 to the preceding relation yields $\sum_{k=0}^{\infty}\alpha_{k}(f(x_{s,k})-f^{\star})<\infty$ almost surely; combined with $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, this fact implies there exists a subsequence $\{x_{s,k_{\tau}}\}_{\tau\in\mathbb{N}_{0}}$ such that $\lim_{\tau\rightarrow\infty}f(x_{s,k_{\tau}})=f^{\star}$ with probability $1$. By the continuity of $f(\cdot)$ and the compactness of $\mathcal{X}$, we conclude that the sequence $\{x_{s,k_{\tau}}\}$ converges to a solution of Problem (1).
In addition, by Lemma 14 applied on Eq. (21), there exists a nonnegative random variable $V^{\star}$ such that $\{V_{k}\}\rightarrow V^{\star}$ with probability $1$. Combined with Theorem 27, this implies that the sequence $\{\|x_{s,k}-x^{\star}\|^{2}\}$ converges almost surely. Since $\{x_{s,k_{\tau}}\}_{\tau\in\mathbb{N}_{0}}$ converges almost surely to random solution of Problem (1), we conclude that $\{x_{s,k}\}$ converges to a solution of Problem (1) with probability $1$. ∎
A.2 Additional numerical results for non-iid data
A.2.1 Quadratic problem
Figure 4: Distance to optimality of the global server variable over time for a toy quadratic problem. The stepsize values for all methods are listed in the legend.
To demonstrate the robustness of AREA to heterogeneous local objectives, we construct a toy one-dimensional quadratic problem by defining $f_{i}(x)=\frac{1}{2}(100ix-1)^{2}$ for $i\in[50]$. The rates $\lambda_{i}$ for $i\in[50]$ were sampled from $\mathcal{N}(10,3)$, and we set $\Delta=4$ for AREA, AS-FedAvg, FedBuff, and MIFA and $\Delta=50$ for S-FedAvg to achieve exact convergence. [^4] We note that S-FedAvg with $\Delta<n$ converges faster but not to the correct optimum. All methods utilized the true gradients $\nabla f_{i}(\cdot)$ (one gradient step per client). The stepsizes for all methods where tuned by using grid search in the range $[10^{-6},10^{-3}]$, and the values yielding the fastest convergence were selected for the plots. The average, minimum, and maximum normalized distances $\|x_{s,t}-x^{\star}\|^{2}/\|x^{\star}\|^{2}$ over time $t$ between the global server variable $x_{s,t}$ and the minimum $x^{\star}$ of Problem (1) over $3,000$ repetitions of the experiment can be seen in Fig. 4. We observe that AS-FedAvg and FedBuff are unable to achieve exact convergence when $f_{i}\neq f_{j}$ for $i\neq j\in[n]$, even when the true gradients are available; in addition, AREA achieves exact convergence faster than S-FedAvg because of its asynchronous nature, while MIFA performs in between.
A.2.2 Cross-silo FL
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
Figure 5: Results for $n=4$ clients and one local SGD step per client. Loss function value and test accuracy are plotted for uniform (left) and non-uniform client update frequencies (right) over time. The stepsize values for all methods are listed in the legend.
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
Figure 6: Results for $n=4$ clients and $50$ local SGD steps per client. Loss function value and test accuracy are plotted for uniform (left) and non-uniform client update frequencies (right) over time. The stepsize values for all methods are listed in the legend.
Our results for the execution of a single local SGD step at each client and network size $n=4$ are shown in Fig. 5. In Fig. 5 we observe that FedBuff and MIFA outperform AREA and the other methods for both uniform and non-uniform client update frequencies. Comparing the performance of AREA for $n=4$ and $n=128$, we conclude that AREA has better scaling properties relative to the other methods and is better suited for the cross-device setting. Similarly to the $n=128$ case, we also tested all methods with $50$ local SGD steps at each client (Fig. 6). In Fig. 6 we observe that multiple local SGD steps destabilize FedBuff and MIFA. AREA outperforms the other methods with respect to testing accuracy, followed by S-FedAvg.
A.3 Numerical results for iid data
In this subsection, we present our numerical results for the multinomial logistic regression problem in Section 5.1 and iid local data. Like in Section 5.1, we provide results for both the cross-device and the cross-silo FL settings.
A.3.1 Cross-device FL
Our numerical results for $n=128$ and $1$ and $50$ local SGD steps at each client are shown in Figs. 7 and 8, respectively. In Fig. 7 we observe that AREA and FedBuff outperform the remaining methods, with AREA being more numerically stable and yielding tighter confidence bounds. In Fig. 8, we observe that the perfomance of multiple local SGD steps speeds up significantly the convergence of AS-FedAvg. On the other hand, while AREA achieves the second fastest convergence, it also converges to a large error neighborhood of the optimal loss function value due to the large value of the stepsize.
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
Figure 7: Results for $n=128$ clients and one local SGD step per client (iid data). Loss function value and test accuracy are plotted for uniform (left) and non-uniform client update frequencies (right) over time. The stepsize values for all methods are listed in the legend.
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
Figure 8: Results for $n=128$ clients and $50$ local SGD steps per client (iid data). Loss function value and test accuracy are plotted for uniform (left) and non-uniform client update frequencies (right) over time. The stepsize values for all methods are listed in the legend.
A.3.2 Cross-silo FL
Our numerical results for $n=4$ and $1$ and $50$ local SGD steps at each client are shown in Figs. 9 and 10, respectively. In Fig. 9, we observe that all methods achieve comparable performance with the exception of S-FedAvg under non-uniform client update frequencies. Conversely, in Fig. 10, we observe that the execution of multiple serial SGD steps destabilizes MIFA and FedBuff, and as in the cross-device setting, boosts the convergence of AS-FedAvg.
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
Figure 9: Results for $n=4$ clients and one local SGD step per client (iid data). Loss function value and test accuracy are plotted for uniform (left) and non-uniform client update frequencies (right) over time. The stepsize values for all methods are listed in the legend.
(a) Loss, uniform client updates
(b) Test accuracy, uniform client updates
(c) Loss, non-uniform client updates
(d) Test accuracy, non-uniform client updates
Figure 10: Results for $n=4$ clients and $50$ local SGD steps per client (iid data). Loss function value and test accuracy are plotted for uniform (left) and non-uniform client update frequencies (right) over time. The stepsize values for all methods are listed in the legend.